### v1

# Almost Complete Intersections: Regularity of powers and Rees algebra

Abstract.
An almost complete intersection ideal can be seen as a $d$-sequence ideal with the minimal number of generators being one more than its height. In this paper, we give exact formulas for the regularity of powers of graded almost complete intersection ideals. We also present the minimal bigraded free resolutions of Rees algebras associated with linear type ideals having one generator more than their grade and study the properties of Cohen-Macaualyness, Koszulness associated to their diagonals.
Key words and phrases:
Almost Complete Intersections, Regularity, Koszul algebra, Cohen-Macaulay, Diagonal Subalgebra
2020 Mathematics Subject Classification:
Primary 13D02, 13J30,13C05; Secondary 13H10
Introduction
All rings in this paper are considered to be standard graded $K$-algebras ($K$ being a field), unless otherwise specified. Rees algebras and symmetric algebras are important classes of bigraded algebras which have powers and symmetric powers of ideals respectively as its graded components. Corresponding to a graded ideal, the defining equations of the symmetric algebra can be given by the presentation of the ideal, but it is not easy to obtain the associated free resolution. Whereas in case of Rees algebras, even finding the explicit defining equations are found to be challenging in general. Ideals of linear type are a class of ideals for which the corresponding symmetric algebras and Rees algebras are isomorphic, which is equivalent to the defining equations of Rees algebras being linear in the second set of variables. Complete intersection ideals and $d$-sequence ideals are some classes of ideals of this type. In case of ideals generated by regular sequences, Koszul complexes resolve the ideals and the Eagon-Northcott complex resolves the minimal free resolution of the associated Rees algebra ([11, 12]). In [24], Rees algebras of certain Pfaffians ideals generated by $d$-sequence which come corresponding to alternating matrices are discussed. There are other special cases for which the correponding Rees algebras are studied [11, 33, 34, 28, 14, 26].
Another class of ideals of linear type are almost complete intersections. An ideal $I$ in a standard graded $K$-algebra $A$ is said to be an almost complete intersection if the minimal number of generators of $I$ is one more than the height of $I$ along with the property that for all minimal primes $\mathfrak{p}$ of $I$ in $A$, $I_{\mathfrak{p}}$ is a complete intersection. It is proved in [21] that these ideals are contained in the class of $d$-sequence ideals. Further some study has been done on their regularity [20, 31], combinatorial characterization of graphs which correspond to these ideals ([21]) and their corresponding Rees algebras [13, 21]. These classes of ideals can be seen as losing the complete intersection property by one generator, yet unlike the ideals generated by regular sequences, little is known in generality regarding the homological invariants associated to them.
Minimal free resolutions of ideals are one of the main tools which helps in studying their structure. Finding these resolutions corresponding to arbitrary ideals in general, is a difficult problem. Applying stronger conditions on the ideal helps in a better understanding of the homological invariants associated to them. In Section 3 (Proposition 3.1) we give the minimal bigraded free resolution of the Rees algebra corresponding to an ideal satisfying the following properties: $(1)$ ideal is of linear type, $(2)$ minimal number of generators of the ideal is one more than its grade. This is a generalization of a result in [5] extending it to the case of non-equigenerated ideals. As a consequence, one obtains resolutions of the bigraded algebras associated to almost complete intersection ideals. In particular, it shows that in order to obtain the minimal free resolution of the Rees algebra of almost complete intersection ideals, it suffices to have information about the resolution of $I$.
A significant algebraic invariant which measures the complexity of a module is Castelnuovo-Mumford regularity. It is proved in [10, 22] that for every homogeneous ideal $I$, the regularity of $I^{s}$ is of the form $as+b$ for $s\gg 0$ and non-negative constants $a,b$. While computing the value of $b$ is often considered to be a challenging problem, the value of $a$ is widely understood in literature. In Section 2 (Theorem 2.3), we derive the constant $b$ for powers of almost complete intersections thereby extending the result of Shen and Zhu [31, Theorem 3.1] to include the case when the ideal is non-equigenerated.
Let $c,e\geq 0$ be two integers with $(c,e)\neq(0,0)$. Then the diagonal subalgebra of a bigraded algebra $R$ along the $(c,e)$-diagonal $\Delta=\{(cs,es):s\in\mathbb{Z}\}$ is defined as the graded algebra $R_{\Delta}:=\bigoplus_{i\geq 0}R_{(ci,ei)}$ (cf. [32]). From a geometric point of view, they are the homogeneous coordinate rings of embeddings of blow-ups of projective varieties along a subvariety. For $c,e\geq 0$, and a homogeneous ideal $I$ in a graded ring $A$, $K[(I^{e})_{c}]$, the $K$-algebra generated by the $c^{th}$ Veronese subring of $I^{e}$, can be identified as a $(c,e)$-diagonal of the Rees algebra if $c\geq ed+1$, thereby identifying them as the subalgebras of the Rees algebras in a natural way [32, 9]. This association facilitates an algebraic approach to studying $K[(I^{e})_{c}]$.
A standard graded $K$-algebra is Koszul if the minimal free resolution of the corresponding residue field $K$ has a linear resolution. Some articles that have discussed these properties of the diagonal subalgebras of Rees algebras of ideals are [32, 9, 6, 23, 1] which includes the cases of complete intersections, residual intersections etc. In Section 4, we study the ring-theoretic properties like Cohen-Macaulayness, Koszulness of the diagonals of Rees algebras associated to certain specific classes of ideals of linear type. For a complete intersection ideal $I$ in a polynomial ring minimally generated by $r$ elements of degree $d$, it is proved in [9, Corollary 6.10] that for $c\geq\frac{d(r-1)}{r}$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul and an improved bound of $c\geq\frac{d}{2}$ is given for the case of $r=2$, $d=2$ in [6, Corollary 3.3] and for the general case of $r=2$ and any $d$ in [23, Theorem 3.1]). In Proposition 4.1$(1)$, we prove a similar result for an almost complete intersection ideal $I$ minimally generated by $3$ elements in degree $d$ in a graded polynomial ring, which is, $\mathcal{R}(I)_{\Delta}$ is Koszul for $c\geq\frac{d}{3}$ and $e>0$.
1. Preliminaries
Following are some notations used throughout the article.
  1. a)
    $K$ denotes a field of characteristic zero.
  2. b)
    For an ideal $I$, $\mu(I)$ denotes the minimal number of generators of $I$ and ht$(I)$ denotes the height of $I$.
  3. c)
    $A$ denotes a standard graded ring of the form $B/I$ where $B$ is the standard graded polynomial ring over the field $K$ and $I$ is an ideal of $B$.
Let $M$ be a finitely generated graded $A$-module which admits the following graded minimal free resolution
$\mathbb{F}:\cdots\stackrel{{\scriptstyle d_{\mathnormal{i+1}}}}{{\longrightarrow}}\bigoplus_{{\mathnormal{j}}}A(-\mathnormal{j})^{{{\beta_{\mathnormal{i},\mathnormal{j}}}}}\stackrel{{\scriptstyle d_{\mathnormal{i}}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle d_{\mathnormal{2}}}}{{\longrightarrow}}\bigoplus_{{\mathnormal{j}}}A(-\mathnormal{j})^{{{\beta_{\mathnormal{1},\mathnormal{j}}}}}\stackrel{{\scriptstyle d_{\mathnormal{1}}}}{{\longrightarrow}}\bigoplus_{{\mathnormal{j}}}A(-\mathnormal{j})^{{{\beta_{\mathnormal{0},\mathnormal{j}}}}}\longrightarrow 0.$
Set $t_{\mathnormal{i}}^{A}(M):=\sup\{\mathnormal{j}:\beta_{\mathnormal{i},\mathnormal{j}}^{A}(M)\neq 0\}$ where $\beta_{i,j}$ is the $(i,j)^{\text{th}}$ Betti number of $M$ and by convention, $t_{\mathnormal{i}}^{A}(M)=-\infty$ if $F_{\mathnormal{i}}=0.$ Castelnuovo-Mumford regularity is then defined as follows:
$\operatorname{reg}_{A}(M)=sup\{j-i:\beta_{\mathnormal{i},\mathnormal{j}}^{A}(M)\neq 0;\,i,j\in\mathbb{N}\}=sup\{t_{i}^{A}(M)-i:i\in\mathbb{N}\}.$
Castelnouvo-Mumford regularity measures the complexity of the minimal free resolution of a finitely generated $A$-module $M$. We use the notation $\operatorname{reg}(M)$ when the ring $A$ is evident.
Similarly, considering the bigraded minimal free resolution of a finitely generated bigraded module $N$ over a bigraded ring $R$, one obtains biregularities. They are defined as follows,
  1. (1)
    $\text{reg}_{y}(N)=\text{sup}\{b\in\mathbb{Z}:\,\beta_{i,(a,b+i)}\neq 0,\text{ for some }i,a\in\mathbb{Z}\}$ denotes the $y$-regularity of $N$.
  2. (2)
    $\text{reg}_{x}(N)=\text{sup}\{a\in\mathbb{Z}:\,\beta_{i,(a+i,b)}\neq 0,\text{ for some }i,b\in\mathbb{Z}\}$ denotes the $x$-regularity of $N$.
The following lemma gives an inequality for the regularity of modules in a short exact sequence.
Lemma 1.1 (Regularity Lemma).
[17, Lemma 3.1] Let $0\rightarrow M\rightarrow N\rightarrow P\rightarrow 0$ be a short exact sequence of finitely generated graded $A$-modules. Then, the following holds.
  1. (1)
    If $\operatorname{reg}(M)\neq\operatorname{reg}(P)+1,$ then $\operatorname{reg}(N)=\max\{\operatorname{reg}(M),\operatorname{reg}(P)\}$.
  2. (2)
    If $\operatorname{reg}(N)\neq\operatorname{reg}(P),$ then $\operatorname{reg}(M)=\max\{\operatorname{reg}(N),\operatorname{reg}(P)+1\}$.
  3. (3)
    If $\operatorname{reg}(M)\neq\operatorname{reg}(N),$ then $\operatorname{reg}(P)=\max\{\operatorname{reg}(M)-1,\operatorname{reg}(N)\}$.
Similar to the regularity lemma, one can obtain a lower bound for the depth of a module in terms of that of the free modules in the graded resolution of the same.
Lemma 1.2 (Depth Lemma).
Let $M$ be a finitely generated $A$-module, with the free modules in its graded minimal free resolution denoted by $F_{i}$, $i=0,\ldots,\operatorname{proj\,dim}_{A}M$. Then the following inequality holds,
$\operatorname{depth}\,(M)\geq\min\{\operatorname{depth}\,(F_{i})-i;\,i\geq 0\}.$
Let $\{\bf{a}\}$=$=\{a_{1},\ldots,a_{n}\}$ be a sequence, $I=\langle a_{1},\ldots,a_{n}\rangle$ and for $i=1,\ldots,n$, $I_{i}=\langle a_{1},\ldots,a_{i}\rangle$. Then $\{\bf{a}\}$ is said to be a $d$-sequence with respect to a module $M$ if it satisfies the condition $(I_{i-1}M:_{M}a_{i})\cap IM=I_{i-1}M$ for $i=1,\ldots,n$.
Note 1.3.
Following are some properties of $d$-sequences and regular sequences which will be used in this article.
  1. (1)
    If $\{f_{1},\ldots,f_{n}\}$ forms a $d$-sequence in a ring $A$, then
    1. (a)
      $(\langle f_{1},\ldots,f_{n-1}\rangle:f_{n}^{s})=(\langle f_{1},\ldots,f_{n-1}\rangle:f_{n})$ for all $s\geq 1$.
    2. (b)
      For $s\geq 1$, ${(\langle f_{1},\ldots,f_{i-1}\rangle+I^{s}:f_{i}})=(\langle f_{1},\ldots,f_{i-1}\rangle:f_{i})+I^{s-1}$ for all $1\leq i\leq n$ ([31, Observation 2.4]).
  2. (2)
    If $\{f_{1},\ldots,f_{n}\}$ forms a regular sequence, then $(\langle f_{1},\ldots,f_{i-1}\rangle:f_{i})=\langle f_{1},\ldots,f_{i-1}\rangle$ for all $1\leq i\leq n$.
For a field $K$, $A$ is said to be a standard graded $K$-algebra if $A$ can be identified as $B/I$, where $B$ is a graded polynomial ring over $K$ ($K$-algebra) with the degree of the variable being $1$ and $I$ is its homogeneous ideal. For an equigenerated ideal $\mathfrak{I}$ of $A$, the Rees algebra of $\mathfrak{I}$ in $A$ is a bigraded $K$-algebra defined as $\mathcal{R}(\mathfrak{I})=\bigoplus_{n\geq 0}{\mathfrak{I}}^{n}$.
Let $c,e\geq 0$ be two integers with $(c,e)\neq(0,0)$. Then the diagonal subalgebra of a bigraded algebra $R$ along the $(c,e)$-diagonal $\Delta=\{(cs,es):s\in\mathbb{Z}\}$ is defined as the graded algebra $R_{\Delta}=\bigoplus_{i\geq 0}R_{(ci,ei)}$ (cf. [32]).
Remark 1.4.
Following are some results related to the Koszulness and Cohen-Macaulayness of the diagonals of bigraded algebras which will be used in the paper.
  1. (1)
    For $\Delta=(c,e)$ and a bigraded $K$-algebra $R$ with the $i^{th}$ Betti number corresponding to the shift $(a,b)$ being denoted by $\beta_{i,a,b}$, $R_{\Delta}$ is Koszul if $\max\{\dfrac{a}{c},\dfrac{b}{e}:\beta_{i,a,b}\neq 0\}\leq i+1$ for every $i$ ([9, Theorem 6.2]).
  2. (2)
    Let $I=\langle f_{1},\ldots,f_{s}\rangle$ be a homogeneous ideal of a standard graded $K$-polynomial ring $B=K[x_{1},\ldots,x_{m}]$ with deg$(f_{j})=d_{j}$. Let $d=\text{max}\{d_{1},\ldots,d_{s}\}$ and $u=\sum_{i=1}^{s}d_{i}$. If $\mathcal{R}(I)$ is Cohen-Macaulay, then for $\alpha=\text{min}\{(e-1)d+u-m,e(u-m)\}$ and $\beta=\text{min}\{(e-1)d+u-d_{1},e(u-d_{1})\}$, $\mathcal{R}(I)_{\Delta}$ is Cohen-Macaulay if $c>\text{max}\{\alpha,\beta,de\}$ [27, Theorem 4.5].
2. Regularity of powers of Almost Complete Intersections
In [4], the authors give the regularity of powers of an equigenerated complete intersection $I$ generated by a regular sequence of length $n$ and in degree $d$ as,
$\operatorname{reg}(I^{s})=ds+(d-1)(n-1).$ (1)
Further, in [31], Shen and Zhu give exact formulas for the regularity of powers of equigenerated almost complete intersections.
In this section, we extend their result to give the regularity of powers of almost complete intersections in the non-equigenerated case as well. Our result coincides with the result in [31] when the ideal is equigenerated. Note that an upper bound for the same is given in [20].
Setup 1.
Let $A$ be a standard graded $K$-algebra and $I$ be a homogeneous ideal of $A$ generated by a $d$-sequence $\{f_{1},\ldots,f_{n}\}$ in degrees $d_{1},\ldots,d_{n}$ with $\{f_{1},\ldots,f_{n-1}\}$ being a regular sequence and $J=\langle f_{1},\ldots,f_{n-1}\rangle$. Assume that $\operatorname{reg}(A/I)<\sum_{l=1}^{n-1}d_{l}-n+1$ and $d_{n}=\max\{d_{i}\mid i=1,\ldots,n\}$.
Lemma 2.1.
Consider the notation as assumptions as in Setup 1. Then,
$\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{2}}\right)=\sum_{l=1}^{n}d_{l}-n.$
Proof.
Consider the short exact sequence,
$0\longrightarrow\dfrac{A}{J:f_{n}}(-d_{n})\stackrel{{\scriptstyle.f_{n}}}{{\longrightarrow}}\dfrac{A}{J}\longrightarrow\dfrac{A}{I}\longrightarrow 0$
Here, $\operatorname{reg}\left(\dfrac{A}{J}\right)=\sum_{l=1}^{n-1}d_{l}-(n-1)>\operatorname{reg}\left(\dfrac{A}{I}\right)$. Hence,
$\operatorname{reg}\left(\dfrac{A}{J:f_{n}}\right)=\sum_{l=1}^{n-1}d_{l}-(n-1)-d_{n}.$ (2)
The result is now proved by descending induction on $i$. Consider the following short exact sequence for looking at the base case $i=n-1$ of descending induction.
$0\longrightarrow\dfrac{A}{(J:f_{n}^{2})}(-2d_{n})\stackrel{{\scriptstyle.f_{n}^{2}}}{{\longrightarrow}}\dfrac{A}{J}\longrightarrow\dfrac{A}{J+\langle f_{n}^{2}\rangle}\longrightarrow 0$
As a consequence of Note 1.3, $\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}^{2}\rangle}\right)=\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}\rangle}\right)=\sum_{l=1}^{n-1}d_{l}-n+1-d_{n}.$ Then $\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}^{2}\rangle}(-2d_{n})\right)=\operatorname{reg}\left(\dfrac{A}{J}\right)+d_{n}$.
Therefore, from Lemma 1.1$(3)$, $\operatorname{reg}\left(\dfrac{A}{J+\langle f_{n}^{2}\rangle}\right)=\sum_{l=1}^{n}d_{l}-n$. Now, assume that $i\leq n-2$ and the statement holds for $i$. Thus, considering the short exact sequence,
$0\longrightarrow\dfrac{A}{(\langle f_{1},\ldots,f_{i-1}\rangle+I^{2}):f_{i}}(-d_{i})\stackrel{{\scriptstyle.f_{i}}}{{\longrightarrow}}\dfrac{A}{\langle f_{1},\ldots,f_{i-1}\rangle+I^{2}}\longrightarrow\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{2}}\longrightarrow 0$
Since $\{f_{1},\ldots,f_{n}\}$ forms a $d$-sequence with $\{f_{1},\ldots,f_{n-1}\}$ forming a regular sequence, from Note 1.3,
$\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i-1}\rangle+I^{2}):f_{i}}\right)(-d_{i})=\operatorname{reg}\left(\dfrac{A}{I}\right)+d_{i}<\sum_{l=1}^{n-1}d_{i}-(n-1)+d_{i}\leq\sum_{l=1}^{n}d_{i}-(n-1)\leq$
$\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{2}}\right)+1$. Hence from Lemma $(1)$,
$\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i-1}\rangle+I^{2}}\right)=\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{2}}\right)=\sum_{l=1}^{n}d_{l}-n.$
Lemma 2.2.
Consider the notation and assumptions as in Setup 1. Then, for $s\geq 2$ and $i=0,\ldots,n-1$,
$\operatorname{reg}\left(\dfrac{A}{(J:f_{n})+\langle f_{n}^{s}\rangle}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}.$
Proof.
First we claim that,
$\operatorname{reg}\left(\dfrac{A}{(J:f_{n})+\langle f_{n}^{s}\rangle}\right)=\operatorname{reg}\left(\dfrac{A}{J:f_{n}}\right)+sd_{n}-1.$
For this consider the following short exact sequence.
$0\longrightarrow\dfrac{A}{J:f_{n}^{s+1}}(-sd_{n})\stackrel{{\scriptstyle.f_{n}^{s}}}{{\longrightarrow}}\dfrac{A}{J:f_{n}}\longrightarrow\dfrac{A}{(J:f_{n})+\langle f_{n}^{s}\rangle}\longrightarrow 0$
From Note 1.3, one obtains $\operatorname{reg}\left(\dfrac{A}{J:f_{n}^{s+1}}\right)(-sd_{n})=\operatorname{reg}\left(\dfrac{A}{J:f_{n}}\right)+sd_{n}$. Then from Lemma 1.1$(3)$,
$\operatorname{reg}\left(\dfrac{A}{(J:f_{n})+\langle f_{n}^{s}\rangle}\right)=\operatorname{reg}\left(\dfrac{A}{J:f_{n}}\right)+sd_{n}-1.$ (3)
Now substituting the value of $\operatorname{reg}\left(\dfrac{A}{J:f_{n}}\right)$ from Equation (2) in Equation (3), we get the required result.
As a consequence of the above results, one can obtain the regularity of powers of almost complete intersections.
Theorem 2.3.
Consider the notation and assumptions as in Setup 1. Then for all $s\geq 2$ and $i=0,\ldots,n-1$, the following is true,
$\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{s}}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}.$
Proof.
We prove by induction on $s\geq 2$. The base case for $s=2$ is a consequence of Lemma 2.1. Let $s\geq 3$. For all $j,s\in\mathbb{N}$ with $j<n$, we look at the following short exact sequence.
$0\longrightarrow\dfrac{A}{(\langle f_{1},\ldots,f_{i}\rangle+I^{s}):f_{i+1}}(-d_{i+1})\stackrel{{\scriptstyle.f_{i+1}}}{{\longrightarrow}}\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{s}}\longrightarrow\dfrac{A}{\langle f_{1},\ldots,f_{i+1}\rangle+I^{s}}\longrightarrow 0$ (4)
We prove the result by descending induction on $i$. Consider the base case of $i=n-1$ of the descending induction in Equation (4). Then clearly, $\langle f_{1},\ldots,f_{(n-1)+1}\rangle+I^{s}=I$. From Note 1.3 and Lemma 2.2,
$\operatorname{reg}\left(\dfrac{A}{(J+I^{s}):f_{n}}(-d_{n})\right)=\operatorname{reg}\left(\dfrac{A}{(J:f_{n})+I^{s-1}}\right)+d_{n}=\sum_{l=1}^{n}d_{l}-n+(s-3)d_{n}+d_{n}=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}\geq\sum_{l=1}^{n-1}d_{l}-(n-1)+1>\operatorname{reg}\left(\dfrac{A}{I}\right)+1$. Hence $\operatorname{reg}\left(\dfrac{A}{J+I^{s}}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}$ by Lemma $(1)$.
Now assume $i\leq n-2$ and that the result holds for $i+1$. In the short exact sequence (4), by induction hypothesis on $i$, $\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i+1}\rangle+I^{s}}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}$. Since $\{f_{1},\ldots,f_{j}\}$ forms a $d$-sequence for $1\leq j\leq n$ and a regular sequence for $11\leq j\leq n-$ 1, from Note 1.3, $\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i}\rangle+I^{s}):f_{i+1}}\right)=\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{s-1}}\right)$. This by induction hypothesis on $s$ is equal to $\sum_{l=1}^{n}d_{l}-n+(s-3)d_{n}$.
Now, the final result is a consequence of the inequality
$\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i}\rangle+I^{s}):f_{i+1}}\right)+d_{i+1}<\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i+1}\rangle+I^{s}}\right)+1$
and Lemma $(1)$. ∎
Note 2.4.
The above theorem fails if $d_{n}$ is not the maximal degree generator of $I$. For example consider $I=\langle f_{1},f_{2},f_{3},f_{4}\rangle=\langle x_{26}x_{45}-x_{25}x_{46}+x_{24}x_{56},x_{16}x_{45}-x_{15}x_{46}+x_{14}x_{56},x_{16}x_{25}x_{34}-x_{15}x_{26}x_{34}-x_{16}x_{24}x_{35}+x_{14}x_{26}x_{35}+x_{15}x_{24}x_{36}-x_{14}x_{25}x_{36},x_{36}x_{45}-x_{35}x_{46}+x_{34}x_{56}\rangle\subset K[x_{14},x_{15},x_{16},x_{24},x_{25},x_{26},x_{34},x_{35},x_{36},x_{45},x_{46},x_{56}]=B$. Then $\{f_{1},f_{2},f_{3},f_{4}\}$ forms a $d$-sequence with $\{f_{1},f_{2},f_{3}\}$ being a regular sequence. But $\operatorname{reg}\left(\dfrac{B}{I^{3}}\right)=8$ which is not equal to $(2+2+2+3)-4+2=7$.
As a consequence of the above results, the following can be observed regarding the regularity of powers of almost complete intersections of grade $3$. To understand the set of generators of these ideals, we look at the notion of Pfaffians.
Given a skew-symmetric matrix $X$, the Pfaffian of $X$ is defined as the square root of the determinant of $X$. For $l\in\mathbb{N}$, let $\operatorname{Pf}_{\bar{l}}(X)$ denotes the Pfaffian of $X$ obtained by removing the $l^{th}$ row and the corresponding column of $X$ which can be seen as an element in $B=K[X]$. In [7], the authors show that almost complete intersections of grade $3$ arise as the ideal generated by certain Pfaffians of skew-symmetric matrices of some particular form (see Setup 3). Following are some properties of these ideals that can be observed.
  1. (1)
    $B/I$ is Cohen-Macaulay. Since $I$ is perfect in $B$, the observation follows from [12, Section 19].
  2. (2)
    For $t\geq 5$ and $I$ an ideal of type $t-3$,
    1. (a)
      $\text{reg}_{B}(B/I)=t-4$.
    2. (b)
      $(B/I)^{(c)}$ is Koszul for $c\geq\dfrac{t}{3}$.
Lemma 2.5.
Assume the notation and hypotheses as in Setup 3. Then,
  1. (1)
    For an odd integer $t\geq 5$, $\{\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X)\}$ is a $d$-sequence generating set of an almost complete intersection ideal of grade $3$ and type $t-3$.
  2. (2)
    For an even integer $t\geq 6$, $\{\operatorname{Pf}_{\overline{12}}(X),\operatorname{Pf}_{\overline{13}}(X),\operatorname{Pf}_{\overline{23}}(X),\operatorname{Pf}(X)\}$ is a $d$-sequence generating set of an almost complete intersection ideal of grade $3$ and type $t-3$.
Proof.
We have $\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X)\}$ forms a regular sequence from [7, C.2 Lemma]. Further, since $J$ is unmixed, from [13, Proposition 1.3], $\{\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X)\}$ forms a $d$-sequence.
Similarly $(2)$ can be seen as a consequence of [7, C.6 Lemma] and [13, Proposition 1.3]. ∎
Corollary 2.6.
Assume the notation and hypotheses as in Setup 3. Then for $m\geq 2$,
  1. (1)
    If $t\geq 5$ is an odd integer such that , then for $i=1,2,3$,
    $\operatorname{reg}\left(\dfrac{B}{\langle f_{1},\ldots,f_{i}\rangle+I^{m}}\right)=\left(\frac{t-1}{2}\right)(m+2)-5.$
  2. (2)
    If $t\geq 6$ is an even integer, then for $i=1,2,3$,
    $\operatorname{reg}\left(\dfrac{B}{\langle f_{1},\ldots,f_{i}\rangle+I^{m}}\right)=\left(\frac{t}{2}\right)(m+2)-7.$
Proof.
The result can be seen as a consequence of Theorem 2.3 and Lemma 2.5
3. Rees algebra of Almost Complete Intersections
This section discusses the resolutions of the Rees algebra of some classes of ideals $I$ of linear type satisfying a relation between their minimal number of generators and their respective heights. The resolution is obtained as the mapping cone between the resolution of $I$ and the Buchsbaum-Rim complex, coming corresponding to the generators of $I$. The idea has been discussed in [5] for constructing resolutions of equigenerated ideals. We try to extend the result to include the non-equigenerated case as well. As a consequence of this, minimal bigarded free resolutions of almost complete intersections are obtained.
The following are the building blocks for constructing the minimal bigraded free resolutions given in Proposition 3.1.
Let $I$ be an ideal of a standard graded ring $A$ generated by $\{f_{1},\ldots,f_{l}\}$ with $\deg(f_{i})=d_{i}$ and grade $I=l-1$. Let $(\mathbb{F.},\phi.)$ denote the graded minimal free resolution of $A/I$. Now let $S=A[y_{1},\ldots,y_{l}]$ be a polynomial ring and $F$ and $G$ be free $S$-modules of rank $l$ and $2$ respectively. Then the Buchsbaum-Rim complex (we denote it by $\mathbb{C.}$) corresponding to the matrix $\psi=\begin{bmatrix}f_{1}&\cdots&f_{l}\\ y_{1}&\cdots&y_{l}\\ \end{bmatrix}$ is defined as follows,
$0\longrightarrow\wedge^{l}F\otimes(\operatorname{Sym}_{l-3}G)^{*}\stackrel{{\scriptstyle\sigma_{l-2}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\sigma_{2}}}{{\longrightarrow}}\wedge^{3}{F}\otimes(\operatorname{Sym}_{0}G)^{*}\stackrel{{\scriptstyle\epsilon}}{{\longrightarrow}}F\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}G.$
where
  1. (1)
    $(\operatorname{Sym}_{i}G)^{*}$ is the dual of the $i^{th}$ symmetric power of $G$.
  2. (2)
    $\epsilon:\wedge^{3}{F}\otimes(\operatorname{Sym}_{0}G)^{*}\rightarrow F$ is the splice morphism in the Buchsbaum-Rim complex defined as follows.
    Let $\{e_{1},\ldots,e_{l}\}$ be a basis of $F$ and $\{g_{1},g_{2}\}$ be a basis of $G$. We have $(\operatorname{Sym}_{0}G)^{*}=S$ and for $k=\{i_{1},i_{2},i_{3};\;i_{1}<i_{2}<i_{3}\}$, let $e_{i}=e_{i_{1}}\wedge e_{i_{2}}\wedge e_{i_{3}}$ denote an element in $\wedge^{3}{F}$. Then,
    $\epsilon(e_{i})=\sum_{m\subset k,\,|m|=2}\text{sgn}(m\subset k)(\text{det}(\psi_{m}))e_{k\backslash m}$
    where $\text{sgn}(m\subset k)$ is the sign of the permutation of $k$ which puts the elements of $m$ in the first $2$ positions and $\psi_{m}$ is the minor of order $2$ with the columns indexed by the elements of $m$ [5].
  3. (3)
    For $22\leq i\leq l-$ 2, the maps $\sigma_{i}$ are defined as follows,
    $\sigma_{i}(\beta\otimes\gamma)=\sum_{s,t}[\psi^{*}(\beta_{s}^{\prime})](\gamma_{t}^{\prime})\cdot\gamma_{t}^{\prime\prime}\otimes\beta_{s}^{\prime\prime}.$
where $\beta\in(\operatorname{Sym}_{i-1}G)^{*}$ with $\beta=\sum_{s}\beta_{s}^{\prime}\otimes\beta_{s}^{\prime\prime}\in(G)^{*}\otimes(\operatorname{Sym}_{i-2}G)^{*}$ and $\gamma\in\wedge^{i+2}{F}$ with $\gamma=\sum_{s}\gamma_{t}^{\prime}\otimes\gamma_{t}^{\prime\prime}\in\wedge^{1}F\otimes\wedge^{i+1}F$.
Next we mention a map of complexes from $\mathbb{C.}$ to $\mathbb{F.}$ mentioned in [5].
Let $\nu$ be a map of complexes from the Koszul complex corresponding to $\phi_{1}$, denoted by (.$(\wedge^{.}A^{l},c.)$, to the Buchsbaum-Rim complex (.$(\mathbb{C.},\sigma.)$, obtained by extending the identity map at degree 00 and degree $1$ such that the diagram commutes. This map $\nu$ helps in defining the required map of complexes $\alpha:\mathbb{C.}\rightarrow\mathbb{F.}$ which has the following form.
For $21\leq i\leq l-$ 2, $\alpha_{i}:\wedge^{i+2}F\otimes(\operatorname{Sym}_{i-1}G)^{*}\rightarrow F_{i+1}\otimes S$ is given as the composition of the following two maps.
$\alpha_{i}:\wedge^{i+2}F\otimes(\operatorname{Sym}_{i-1}G)^{*}\stackrel{{\scriptstyle p\otimes\tau}}{{\longrightarrow}}\wedge^{1}F\otimes\wedge^{i+1}F\stackrel{{\scriptstyle\tiny{\begin{bmatrix}y_{1}&\cdots&y_{l}\end{bmatrix}}\otimes\nu_{i+1}}}{{\longrightarrow}}S\otimes F_{i+1}$ (5)
where $p:\wedge^{i+2}F\rightarrow\wedge^{1}F\otimes\wedge^{i+1}F$ is the co-multiplication map defined as $e_{j_{1}}\wedge e_{j_{2}}\wedge\cdots\wedge e_{j_{i+2}}\mapsto\sum_{i}e_{j_{i}}\otimes e_{j_{1}}\wedge\ldots\wedge\hat{e_{j_{i}}}\wedge\ldots\wedge e_{j_{i+2}}$ and $\tau:(\operatorname{Sym}_{i-1}G)^{*}\rightarrow S$ defined as $g_{1}^{*(i-1-j)}\otimes g_{2}^{*(j)}\mapsto\left\{\begin{array}[]{ccc}1&\text{if}&j=0\\ 0&\text{if}&j\geq 1.\\ \end{array}\right.$
With the notations set, we move on to give the main result in this section.
Proposition 3.1.
Let $B=K[x_{1},\ldots,x_{m}]$ be a standard graded polynomial $K$-algebra and $I=\langle f_{1},\ldots,f_{n+1}\rangle$ be a graded ideal of $B$ of linear type with deg $f_{j}=d_{j}$ for $j=1,\ldots,n+1$ and grade$(I)=n$ (ht$(I)=n$). Let $S=K[x_{1},\ldots,x_{m},y_{1},\ldots,y_{n+1}]$ be a bigraded $K$-algebra with deg $x_{i}=(1,0)$ and deg $y_{j}=(d_{j},1)$ for $i=1,\ldots,m$ and $j=1,\ldots,n+1$. Then the minimal non-standard bigraded free resolution of $\mathcal{R}(I)$ has the form,
$\cdots\stackrel{{\scriptstyle\delta_{4}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 3}S(-j,-1)^{\beta_{4j}}\\ \oplus\\ C_{2}\end{matrix}\stackrel{{\scriptstyle\delta_{3}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 2}S(-j,-1)^{\beta_{3j}}\\ \oplus\\ C_{1}\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 1}S(-j,-1)^{\beta_{2j}}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0.$ (6)
where $\beta_{ij}$ are the $(i,j)^{th}$ Betti numbers of $B/I$ and for $i=1,\dots,n-1$,
$C_{i}=\bigoplus_{j=0}^{i-1}\bigoplus_{1\leq j_{1}<j_{2}<\cdots<j_{i+2}\leq n+1}S(-(d_{j_{1}}+d_{j_{2}}+\cdots+d_{j_{i+2}}),-j-2)$
are the modules which occur in the Buchsbaum-Rim complex $\mathbb{C}.$ corresponding to the matrix $\begin{bmatrix}f_{1}&\cdots&f_{n+1}\\ y_{1}&\cdots&y_{n+1}\\ \end{bmatrix}$ with differentials denoted by $\sigma_{i}$. Let $\alpha:\mathbb{C}.\rightarrow\mathbb{F}.$ be the map of complexes given in Equation (5). Then for all $i=1,\ldots,\operatorname{proj\,dim}_{S}\mathcal{R}(I)$, the differentials $\delta_{i}$ are of the following form.
$\delta_{1}=\begin{bmatrix}y_{1}&\ldots&y_{n+1}\end{bmatrix}\circ\phi_{2},\;\;\delta_{2}=\begin{pmatrix}\alpha_{1}&\phi_{3}\end{pmatrix},\;\;\delta_{i}=\begin{pmatrix}\phi_{i+1}&\alpha_{i-1}\\ O&\sigma_{i-1}\end{pmatrix}.$
Proof.
Let $\mathbb{F}.$ be a graded minimal free resolution of $B/I$ of the following form,
$\cdots\longrightarrow\bigoplus_{j\geq 3}B(-j)^{\beta_{3j}}\stackrel{{\scriptstyle\phi_{3}}}{{\longrightarrow}}\bigoplus_{j\geq 2}B(-j)^{\beta_{2j}}\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\bigoplus_{j=1}^{n+1}B(-d_{j})\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}B\longrightarrow 0,$ (7)
and $\mathbb{C}.$ be the Buchsbaum-Rim complex corresponding to the map $\psi:S^{n+1}\rightarrow S^{2}$ given by the matrix $\begin{bmatrix}f_{1}&\cdots&f_{n+1}\\ y_{1}&\cdots&y_{n+1}\\ \end{bmatrix}$ denoted as follows,
$0\longrightarrow\wedge^{n+1}{S^{n+1}}\otimes(\operatorname{Sym}_{n-2}S^{2})^{*}\stackrel{{\scriptstyle\sigma_{n-1}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\sigma_{2}}}{{\longrightarrow}}\wedge^{3}{S^{n+1}}\otimes(\operatorname{Sym}_{0}S^{2})^{*}\stackrel{{\scriptstyle\epsilon}}{{\longrightarrow}}S^{n+1}\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}S^{2}.$ (8)
Here for $i=1,\ldots,n+1$, $\psi^{*}(g_{1}^{*})(e_{i})=f_{i}$ and $\psi^{*}(g_{2}^{*})(e_{i})=y_{i}$ which determines the entries in the matrices corresponding to the differentials $\sigma_{i}^{\prime}s$. Then the way the differentials are defined in Complex (8), one obtains
$C_{i}=\bigoplus_{j=0}^{i-1}\bigoplus_{1\leq j_{1}<j_{2}<\cdots<j_{i+2}\leq n+1}S(-(d_{j_{1}}+d_{j_{2}}+\cdots+d_{j_{i+2}}),-j-2)$
for $i=1,\ldots,n-1$.
Consider (.$(\wedge^{.}B^{n+1},k.)$ to be the Koszul complex corresponding to $\phi_{1}$. Now since for $\{j_{1},j_{2}\}\subseteq[n+1]$ such that $j_{1}\neq j_{2}$, $k_{2}(e_{j_{1}}\wedge e_{j_{2}})=f_{j_{1}}e_{j_{2}}-f_{j_{2}}e_{j_{1}}\in$ Ker ϕ1=subscriptitalic-$\phi_{1}=$ Im $\phi_{2}$, there exists $\beta_{j_{1}j_{2}}\in F_{2}$ such that $\phi_{2}(\beta_{j_{1}j_{2}})=i(k_{2}(e_{j_{1}}\wedge e_{j_{2}}))$ where $i$ is the natural map from $\wedge^{0}B^{n+1}$ to $F_{1}$. Let $\alpha:\mathbb{C.}\rightarrow\mathbb{F.}$ be the map of complexes. Then $\alpha_{1}:\wedge^{3}{S^{n+1}}\otimes(\operatorname{Sym}_{0}S^{2})^{*}\rightarrow S\otimes F_{2}$ will be defined as, for $\{j_{1},j_{2},j_{3}\}\subseteq[n+1]$, $\alpha_{1}((e_{j_{1}}\wedge e_{j_{2}}\wedge e_{j_{3}})\otimes 1)=y_{j_{1}}\beta_{j_{2}j_{3}}-y_{j_{2}}\beta_{j_{1}j_{3}}+y_{j_{3}}\beta_{j_{1}j_{2}}.$
With the details above, one obtains the required resolution (6) by considering the mapping cone of $\alpha$ (arguments similar to the ones in [5]). Further, the resolution is minimal since the entries in the matrices given by the differentials $\delta_{i}$ are from $\mathfrak{m}S$, where $\mathfrak{m}$ is the bigraded maximal ideal of $S$. Note that the linear shifts in the second variable is contributed by the differentials in the minimal free resolution of $B/I$ and the higher degree shifts for the same are contributed by $\alpha_{i}^{\prime}s$ and $\sigma_{i}^{\prime}s$. ∎
Note 3.2.
Let $\beta_{i}$ be the $i^{th}$ total Betti number of $B/I$. Then from the way the maps are defined it is clear that $d_{i}$, $\alpha_{i}$, $\delta_{i}$ represents matrices of order $\beta_{i-1}\times\beta_{i}$, $\beta_{i+1}\times i\cdot{{n+1}\choose{i+2}}$ and $(i-1)\cdot{{n+1}\choose{i+1}}\times i\cdot{{n+1}\choose{i+2}}$ respectively.
Remark 3.3.
Note that in Proposition 3.1 if $I$ is considered to be a graded ideal of $B$ with $\mu(I)=\text{ht}(I)+1$ and not necessarily of linear type, then the Complex (6) gives a minimal bigraded free resolution of the $\operatorname{Sym}(I)$.
Remark 3.4.
Let $I$ be an graded ideal of $B$ of linear type such that $\mu(I)=\text{ht}(I)+1$. Then,
$\operatorname{proj\,dim}_{S}\mathcal{R}(I)=\left\{\begin{array}[]{ccc}n&\text{if}&\text{pd }(B/I)=n\\ k-1&\text{if}&\text{pd }(B/I)=k,\,k>n.\\ \end{array}\right.$
Following observations can be made on the nature of the generators of $I$ as a consequence of the explicit minimal bigraded free resolutions of the symmetric algebra of $I$ obtained from Proposition 3.1.
Note that for an ideal $I$ in $B$, it’s generating set $\{a_{1},\ldots,a_{n}\}$ forms a strong $s$-sequence (w.r.t an admissible term order for the monomials in $y_{i}$ with $y_{1}<y_{2}<\ldots<y_{n}$) if for $\operatorname{Sym}(I)\cong S/J$ where $S$ is a bigraded polynomial ring $S=B[y_{1},\ldots,y_{n}]$ and $J$ the defining ideal of the symmetric algebra, $in(J)=\langle L_{1}y_{1},\ldots,L_{n}y_{n}\rangle$ where $L_{i}=(a_{1},\ldots,a_{i-1}:a_{i})$ and $L_{1}\subseteq L_{2}\subseteq\ldots\subseteq L_{n}$ (cf. [16]).
Proposition 3.5.
Let $I$ be graded ideal of a standard graded polynomial ring $B$ satisfying the assumptions in Remark 3.3. Then,
  1. (1)
    $I$ is generated by a strong $s$-sequence.
  2. (2)
    $I$ is generated by a $d$-sequence if $I$ is of linear type.
Proof.
From the minimal bigraded free resolution of the $\operatorname{Sym}(I)$ in Complex (6), it is clear that the $y$-regularity of the $\operatorname{Sym}(I)$ is 00. The result then follows from [30, Corollary 3.2].
Further when the ideal is of linear type, $I$ is generated by a $d$-sequence can be seen as a consequence of [25, Theorem 2.9]. ∎
Some important classes of ideals for which Proposition 3.1 gives the bigraded free resolutions of the corresponding symmetric algebras are the following.
  1. (1)
    If $I$ is a graded ideal of $B$ of the form $I=J+\langle g\rangle$ where $J$ is a complete intersection and $g\in B$ such that $I$ is minimally generated by the generators of the ideal $J$ and $g$.
  2. (2)
    $d$-sequence ideals satisfying the condition that the grade of the ideal is one more than the minimal number of generators of the ideal.
  3. (3)
    Almost complete intersections.
In the remaining part of this section, we give the explicit resolutions of the Rees algebra of some classes of ideals satisfying the hypothesis of Proposition 3.1, in particular almost complete intersections of grade $2$, grade $3$ etc.
3.1. Almost complete intersections of grade $2$
Setup 2.
Let $B$ be a standard graded polynomial ring and $I=\langle f_{1},f_{2},f_{3}\rangle\subseteq B$ be an almost complete intersection of grade $2$ and perfect with deg $f_{i}=d_{i}$, $i=1,2,3$. Further assume that the presentation matrix of $I$ has the property that every non-zero entry in the $i^{th}$ row is of degree $\beta_{i}$.
Proposition 3.6.
Consider the ideal $I$ as in Setup 2 and the notations as in Proposition 3.1. Then the minimal bigraded free resolution of the corresponding $\mathcal{R}(I)$ is of the following form:
$0\longrightarrow\begin{matrix}S(-(d_{1}+d_{2}+d_{3}),-2)\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}S(-(d_{k}+\beta_{k}),-1)\\ \oplus\\ S(-(d_{l}+\beta_{l}),-1)\end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0$
where $\beta_{k}$ and $\beta_{l}$, $k,l\in\mathbb{N}$, corresponds to the degrees of elements from two different columns of the presentation matrix of $I$.
Proof.
From Hilbert-Burch theorem [12, Theorem 20.15], the graded minimal free resolution of $B/I$ has the form,
$0\longrightarrow\begin{matrix}B(-(d_{k}+\beta_{k}))\\ \oplus\\ B(-(d_{l}+\beta_{l}))\end{matrix}\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\begin{matrix}B(-d_{1})\\ \oplus\\ B(-d_{2})\\ \oplus\\ B(-d_{3})\end{matrix}\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}B\longrightarrow 0$
where $\phi_{1}=\left[\begin{matrix}f_{1}&f_{2}&f_{3}\end{matrix}\right]$ and $\phi_{2}=\left[\begin{matrix}z_{11}&z_{12}\\ z_{21}&z_{22}\\ z_{31}&z_{32}\\ \end{matrix}\right]$ and further $\beta_{k}$ and $\beta_{l}$ come corresponding to some non-zero entries $z_{ks}$ and $z_{lt}$ respectively in $\phi_{2}$ such that $s-t\cong 1\mod(2)$.
With a fixed ordered basis, the Buchsbaum-Rim complex ($\mathbb{C}$-complex) has the form,
$0\longrightarrow\begin{matrix}S(-(d_{1}+d_{2}+d_{3}),-2)\\ \end{matrix}\stackrel{{\scriptstyle\epsilon}}{{\longrightarrow}}S^{3}\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}S^{2}$
where $\psi=\begin{bmatrix}f_{1}&f_{2}&f_{3}\\ y_{1}&y_{2}&y_{3}\\ \end{bmatrix}$ and $\epsilon=\begin{bmatrix}f_{2}y_{3}-f_{3}y_{2}\\ f_{3}y_{1}-f_{1}y_{3}\\ f_{1}y_{2}-f_{2}y_{1}\end{bmatrix}.$
Let $z$ be a non-zero divisor in $B$ such that $f_{1}=-z(z_{21}z_{32}-z_{22}z_{31}),\,f_{2}=z(z_{11}z_{32}-z_{12}z_{31}),\text{ and }f_{3}=-z(z_{11}z_{22}-z_{12}z_{21})$, which exists from [12], then $\alpha_{1}=\begin{bmatrix}-z(z_{12}y_{1}+z_{22}y_{2}+z_{32}y_{3})\\ z(z_{11}y_{1}+z_{21}y_{2}+z_{31}y_{3})\end{bmatrix}.$
The result thus follows from Proposition 3.1. ∎
Remark 3.7.
Perfect ideals of grade $2$ are characterized by the Hilbert-Burch theorem [12]. Thus, any $3\times 2$ matrix $X$ with entries from a graded polynomial ring $B$ satisfying the condition that grade of the ideal of the corresponding maximal minors ($I_{2}(X)$) is bounded below by $2$ and, a non-zero divisor $z$ in $B$, corresponds to an ideal $zI_{2}(X)$ which has the form mentioned in Setup 2 and thus the associated Rees algebra is resolved using Proposition 3.6.
3.2. Almost complete intersections of grade $3$
Setup 3.
Let $B=K[X]$ denote a standard graded polynomial ring with $X$ being a skew-symmetric matrix of the form,
$X=\begin{bmatrix}0&0&0&x_{14}&x_{15}&\cdots&x_{1t}\\ 0&0&0&x_{24}&x_{25}&\cdots&x_{2t}\\ 0&0&0&x_{34}&x_{35}&\cdots&x_{3t}\\ -x_{14}&-x_{24}&-x_{34}&0&x_{45}&\cdots&x_{4t}\\ -x_{15}&-x_{25}&-x_{35}&-x_{45}&0&\cdots&x_{5t}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\cdots&\vdots\\ -x_{1t}&-x_{2t}&-x_{3t}&-x_{4t}&-x_{5t}&\cdots&0\\ \end{bmatrix}$
of order $t$. Let $I=\langle f_{1},\ldots,f_{4}\rangle$ with deg $f_{i}=d_{i}$, $i=1,\ldots,4$, be an almost complete intersection ideal of $B$ of grade $3$ and type $t-3$ for $t\in\mathbb{N}$, $t\geq 5$.
Proposition 3.8.
Consider the ideal $I$ as in Setup 3. Then the minimal bigraded free resolution of $\mathcal{R}(I)\cong S/J$ where $S=K[X,Y]$ with $Y=\begin{bmatrix}y_{1},\ldots,y_{4}\\ \end{bmatrix}$ and deg $x_{kl}=(1,0)$ and deg $y_{i}=(d_{i},1)$, $i=1,\ldots,4$ and $J$ is the defining ideal of $\mathcal{R}(I)$, has the following form:
  1. (1)
    When $t=2r+1$, $r\in\mathbb{N}$, $r\geq 2$,
    $0\longrightarrow\begin{matrix}S(-4r+1,-2)\\ \oplus\\ S(-4r+1,-3)\end{matrix}\stackrel{{\scriptstyle\delta_{3}}}{{\longrightarrow}}\begin{matrix}S(-2r,-1)^{t-3}\\ \oplus\\ S(-3r+1,-2)^{3}\\ \oplus\\ S(-3r,-2)\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}S(-2r+1,-1)^{t}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0$
  2. (2)
    When $t=2r$, $r\in\mathbb{N}$, $r\geq 3$,
    $0\longrightarrow\begin{matrix}S(-4r+3,-2)\\ \oplus\\ S(-4r+3,-3)\end{matrix}\stackrel{{\scriptstyle\delta_{3}^{\prime}}}{{\longrightarrow}}\begin{matrix}S(-2r+1,-1)^{t-3}\\ \oplus\\ S(-3r+3,-2)^{3}\\ \oplus\\ S(-3r+2,-2)\end{matrix}\stackrel{{\scriptstyle\delta_{2}^{\prime}}}{{\longrightarrow}}\begin{matrix}S(-2r+2,-1)^{t}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{1}^{\prime}}}{{\longrightarrow}}S\longrightarrow 0$
Proof.
Let $I$ be an almost complete intersection of grade $3$ and type $t-3$. Following can be observed from the results in [7].
  1. (1)
    If $t=2r+1$, $r\geq 2$, then the graded minimal free resolution of $B/I$ is given by
    $0\longrightarrow\begin{matrix}B(-2r)^{t-3}\end{matrix}\stackrel{{\scriptstyle\phi_{3}}}{{\longrightarrow}}\begin{matrix}B(-2r+1)^{t}\end{matrix}\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\begin{matrix}B(-r)^{3}\\ \oplus\\ B(-r+1)\end{matrix}\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}B\longrightarrow 0$
  2. (2)
    If $t=2r$, $r\geq 3$, then the graded minimal free resolution of $B/I$ is given by
    $0\longrightarrow\begin{matrix}B(-2r+1)^{t-3}\end{matrix}\stackrel{{\scriptstyle\phi_{3}}}{{\longrightarrow}}\begin{matrix}B(-2r+2)^{t}\end{matrix}\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\begin{matrix}B(-r)\\ \oplus\\ B(-r+1)^{3}\end{matrix}\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}B\longrightarrow 0$
Let $f_{ij}$ denote the element $f_{i}y_{j}-f_{j}y_{i}$ in $S$. With a fixed ordered basis, the respective $\mathbb{C.}$-complexes have the form,
  1. (1)
    $0\longrightarrow\begin{matrix}S(-4r+1,-2)\\ \oplus\\ S(-4r+1,-3)\\ \end{matrix}\stackrel{{\scriptstyle\sigma_{2}}}{{\longrightarrow}}\begin{matrix}S(-3r+1,-2)^{3}\\ \oplus\\ S(-3r,-2)\\ \end{matrix}\stackrel{{\scriptstyle\epsilon}}{{\longrightarrow}}S^{3}\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}S^{2}$
  2. (2)
    $0\longrightarrow\begin{matrix}S(-4r+3,-2)\\ \oplus\\ S(-4r+3,-3)\\ \end{matrix}\stackrel{{\scriptstyle\sigma_{2}}}{{\longrightarrow}}\begin{matrix}S(-3r+3,-2)^{3}\\ \oplus\\ S(-3r+2,-2)\\ \end{matrix}\stackrel{{\scriptstyle\epsilon}}{{\longrightarrow}}S^{3}\stackrel{{\scriptstyle\psi}}{{\longrightarrow}}S^{2}$
where $\psi=\begin{bmatrix}f_{1}&f_{2}&f_{3}&f_{4}\\ y_{1}&y_{2}&y_{3}&y_{4}\\ \end{bmatrix}$, $\epsilon=\begin{bmatrix}f_{23}&f_{24}&f_{34}&0\\ -f_{13}&-f_{14}&0&f_{34}\\ f_{12}&0&-f_{14}&-f_{24}\\ 0&f_{12}&f_{13}&f_{23}\\ \end{bmatrix}$ and $\sigma_{2}=\begin{bmatrix}-f_{4}&-y_{4}\\ f_{3}&y_{3}\\ -f_{2}&-y_{2}\\ f_{1}&y_{1}\\ \end{bmatrix}.$
Then the bigraded free resolutions of $\mathcal{R}(I)$ are consequences of Proposition 3.1.
Corollary 3.9.
The bi-regularities of $\mathcal{R}(I)$ when $I$ is an almost complete intersection of grade $3$ of type $t-3$ are as follows,
  1. (1)
[…]
  2. (2)
[…]
  3. (3)
[…]
3.3. Restriction on the degree of the generators of the ideal and certain colon ideals
Let $I\subseteq B$ such that ht$(I)=n$ and $\mu(I)=n+1$. Let $I=\langle f_{1},\ldots,f_{n+1}\rangle$ such that $f_{1},\ldots,f_{n}$ be $B$-regular and let $J=\langle f_{1},\ldots,f_{n}\rangle$ with deg $f_{i}=d$ for all $i$.
The $\mathbb{C.}$-complex corresponding to $\psi=\begin{bmatrix}f_{1}&\cdots&f_{n+1}\\ y_{1}&\cdots&y_{n+1}\\ \end{bmatrix}$ has the form,
$0\longrightarrow\bigoplus_{l=2}^{n}S(-(n+1)d,-l)\stackrel{{\scriptstyle\sigma_{n-1}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\sigma_{3}}}{{\longrightarrow}}\begin{matrix}S(-4d,-2)^{n+1\choose 4}\\ \oplus\\ S(-4d,-3)^{n+1\choose 4}\\ \end{matrix}\stackrel{{\scriptstyle\sigma_{2}}}{{\longrightarrow}}\begin{matrix}S(-3d,-2)^{n+1\choose 3}\\ \end{matrix}\stackrel{{\scriptstyle\epsilon}}{{\longrightarrow}}S^{n+1}\stackrel{{\scriptstyle{\psi}}}{{\longrightarrow}}S^{2}.$
Now to get a free resolution of $B/I$, consider the following short exact sequence,
$0\longrightarrow\dfrac{B}{(J:f_{n+1})}(-d)\stackrel{{\scriptstyle.f_{n+1}}}{{\longrightarrow}}\dfrac{B}{J}\longrightarrow\dfrac{B}{I}\longrightarrow 0$ (9)
Since $J$ is generated by a regular sequence, the corresponding Koszul complex resolves $B/J$, that is, the graded minimal free resolution of $B/J$ is given by,
$0\longrightarrow B(-nd)\stackrel{{\scriptstyle\eta_{n}}}{{\longrightarrow}}B(-(n-1)d)^{n}\stackrel{{\scriptstyle\eta_{n-1}}}{{\longrightarrow}}\cdots\longrightarrow B(-2d)^{n\choose 2}\stackrel{{\scriptstyle\eta_{2}}}{{\longrightarrow}}B(-d)^{n}\stackrel{{\scriptstyle\eta_{1}}}{{\longrightarrow}}B\longrightarrow 0$
  1. (a)
    Assume $J^{\prime}=(J:f_{n+1})$ to be an equigenerated perfect ideal of grade $2$ with the degree of the generators being $d^{\prime}$, $\mu(J^{\prime})=t$, $t\in\mathbb{N}$ and $J^{\prime}$ being linearly presented. By Hilbert-Burch theorem, the graded minimal free resolution of $B/J^{\prime}$ has the following form:
    $0\longrightarrow B(-(d^{\prime}+1))^{t-1}\stackrel{{\scriptstyle\zeta_{2}}}{{\longrightarrow}}B(-d^{\prime})^{t}\stackrel{{\scriptstyle\zeta_{1}}}{{\longrightarrow}}B\longrightarrow 0$
    Then the mapping cone of the map, multiplication by $f_{n+1}$ in the short exact sequence (9), gives a graded free resolution of $B/I$ (not necessarily minimal) of the following form:
    $0\longrightarrow B(-nd)\stackrel{{\scriptstyle\phi_{n}}}{{\longrightarrow}}B(-(n-1)d)^{n}\stackrel{{\scriptstyle\phi_{n-1}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\phi_{4}}}{{\longrightarrow}}\begin{matrix}B(-(d^{\prime}+d+1))^{t-1}\\ \oplus\\ B(-3d)^{{n\choose{3}}}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{3}}}{{\longrightarrow}}\begin{matrix}B(-(d^{\prime}+d))^{t}\\ \oplus\\ B(-2d)^{{n\choose 2}}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\begin{matrix}B(-d)^{n+1}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}B\longrightarrow 0$
    As a consequence of Proposition 3.1, the following would be a (non-standard) bigraded free resolution of $\mathcal{R}(I)$,
    $0\longrightarrow\bigoplus_{l=2}^{n}S(-(n+1)d,-l)\stackrel{{\scriptstyle\delta_{n}}}{{\longrightarrow}}\begin{matrix}S(-nd,-1)\\ \oplus\\ \bigoplus_{l=2}^{n-1}S(-nd,-l)^{n+1}\end{matrix}\stackrel{{\scriptstyle\delta_{n-1}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\delta_{3}}}{{\longrightarrow}}\begin{matrix}S(-(d^{\prime}+d+1),-1)^{t-1}\\ \oplus\\ S(-3d,-1)^{{n\choose 3}}\\ \oplus\\ S(-3d,-2)^{n+1\choose 3}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}S(-(d^{\prime}+d),-1)^{t}\\ \oplus\\ S(-2d,-1)^{{n\choose 2}}\\ \end{matrix}\stackrel{{\scriptstyle{\delta_{1}}}}{{\longrightarrow}}S\longrightarrow 0$
    Remark 3.10.
    1. 1)
      The ideal $J^{\prime}$ need not always be equigenerated. For example let $I=\langle f_{1},f_{2},f_{3}\rangle$, an ideal in a polynomial ring $B$, be a complete intersection ideal. Then the defining relations of the corresponding Rees algebra is given by the maximal minors of the matrix $X=\begin{bmatrix}f_{1}&f_{2}&f_{3}\\ y_{1}&y_{2}&y_{3}\end{bmatrix}$. Let the relations be denoted by $\{g_{1},g_{2},g_{3}\}$. Then they satisfy the condition that $\{g_{1},g_{2}\}$ are $S$-regular and $(\langle g_{1},g_{2}\rangle:g_{3})=\langle f_{3},y_{3}\rangle$ ([6, Lemma 3.1]), which is clearly not equigenerated.
    2. 2)
      A class of ideals which satisfies the condition (a) are complete intersections minimally generated by $2$ linear terms. For example for $I=\langle x_{1}x_{2},x_{3}x_{4},x_{2}x_{3}\rangle\subset K[x_{1},x_{2},x_{3},x_{4}]$, then $\{x_{1}x_{2},x_{3}x_{4}\}$ forms a regular sequence and $(\langle x_{1}x_{2},x_{3}x_{4}\rangle:x_{2}x_{3})=\langle x_{1},x_{4}\rangle$ is a grade $2$ perfect ideal. In general, as mentioned in Remark 3.7, all perfect ideals of grade $2$ are characterized by the Hilbert Burch theorem.
  2. (b)
    Let $J^{\prime}=(J:f_{n+1})$ be an equigenerated Gorenstein ideal of grade $3$ with the degree of the generators being $d^{\prime}$, $\mu(J^{\prime})=t$ and the ideal $J^{\prime}$ being linearly presented. Then the graded minimal free resolution of $B/J^{\prime}$ has the following form:
    $0\longrightarrow B(-(2d^{\prime}+1))\stackrel{{\scriptstyle\zeta_{1}^{*}}}{{\longrightarrow}}B(-(d^{\prime}+1))^{t}\stackrel{{\scriptstyle\zeta_{2}}}{{\longrightarrow}}B(-d^{\prime})^{t}\stackrel{{\scriptstyle\zeta_{1}}}{{\longrightarrow}}B\longrightarrow 0$
    As mentioned in (a), by considering the mapping cone of the map, multiplication by $f_{n+1}$ in the short exact sequence (9), one obtains a graded free resolution of $B/I$ (not necessarily minimal) of the following form:
    $0\longrightarrow B(-nd)\stackrel{{\scriptstyle\phi_{n}}}{{\longrightarrow}}\cdots\longrightarrow\begin{matrix}B(-(2d^{\prime}+d+1))\\ \oplus\\ B(-4d)^{{n\choose 4}}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{4}}}{{\longrightarrow}}\begin{matrix}B(-(d^{\prime}+d+1))^{t}\\ \oplus\\ B(-3d)^{{n\choose 3}}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{3}}}{{\longrightarrow}}\begin{matrix}B(-(d^{\prime}+d))^{t}\\ \oplus\\ B(-2d)^{{n\choose 2}}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\begin{matrix}B(-d)^{n+1}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}B\longrightarrow 0$
    From Proposition 3.1, the following would be a bigraded free resolution of $\mathcal{R}(I)$,
    $0\longrightarrow\bigoplus_{l=2}^{n}S(-(n+1)d,-l)\stackrel{{\scriptstyle\delta_{n}}}{{\longrightarrow}}\cdots\longrightarrow\begin{matrix}S(-(2d^{\prime}+d+1),-1)\\ \oplus\\ S(-4d,-1)^{{n\choose 4}}\\ \oplus\\ S(-4d,-2)^{n+1\choose 4}\\ \oplus\\ S(-4d,-3)^{n+1\choose 4}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{3}}}{{\longrightarrow}}\begin{matrix}S(-(d^{\prime}+d+1),-1)^{t}\\ \oplus\\ S(-3d,-1)^{{n\choose 3}}\\ \oplus\\ S(-3d,-2)^{n+1\choose 3}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}S(-(d^{\prime}+d),-1)^{t}\\ \oplus\\ S(-2d,-1)^{{n\choose 2}}\\ \end{matrix}\stackrel{{\scriptstyle{\delta_{1}}}}{{\longrightarrow}}S\longrightarrow 0$
Note 3.11.
In both the cases (a) and (b), $d^{\prime}\leq d$ since $J\subseteq J^{\prime}$. This needs to be kept in mind while studying the properties of diagonals in the next section.
4. Application to diagonal subalgebras of almost complete intersections
By extending the idea of the Segre product of algebras, Simis, Trung, and Valla introduced the concept of diagonal subalgebras in [32]. Finding appropriate conditions on a bigraded algebra $R$ such that some algebraic features of $R$ are inherited by $R_{\Delta}$ is one of the main difficulties in the study of diagonal subalgebras.
[…]
Consider the notations as in Proposition 3.1.
4.1. Koszulness of $\mathcal{R}(I)_{\Delta}$
Let $I=\langle f_{1},\ldots,f_{n+1}\rangle$ be an equigenerated ideal of $B$ satisfying the conditions in Proposition 3.1 with deg $f_{i}=d$. Then it is possible to associate a standard bigraded structure to the corresponding Rees algebra by considering it to be the quotient of a standard bigraded polynomial ring $S$ with deg $x_{i}=(1,0)$ and deg $y_{j}=(0,1)$ for $i=1,\ldots,m$ and $j=1,\ldots,n+1$. From the form of the differentials mentioned in Proposition 3.1, the standard bigraded free resolution of the Rees algebra will take the following form,
$\cdots\stackrel{{\scriptstyle\delta_{4}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 3}S(-j+d,-1)^{\beta_{4j}}\\ \oplus\\ C_{2}^{\prime}\end{matrix}\stackrel{{\scriptstyle\delta_{3}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 2}S(-j+d,-1)^{\beta_{3j}}\\ \oplus\\ C_{1}^{\prime}\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 1}S(-j+d,-1)^{\beta_{2j}}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0.$ (10)
where
$C_{i}^{\prime}=\bigoplus_{j=0}^{i-1}S(-(i-j)d,-j-2)^{{n+1}\choose{i+2}}$
for $i=1,\ldots,n-1$. Having defined standard bigraded structure on $\mathcal{R}(I)$, it makes sense to talk about the Koszulness of the corresponding diagonals, since $\Delta$ being an exact functor, when applied to the standard bigraded resolution of $\mathcal{R}(I)$ gives a standard graded resolution of $\mathcal{R}(I)_{\Delta}$.
It is clear from Complex (10) that the $y$-regularity of $\mathcal{R}(I)$ vanishes, and thus $\mathcal{R}(I)_{\Delta}$ is Koszul for $e>0$ [25, Theorem 2.9]. Thus only the $x$-shifts need to be looked into. Let $i$ denote the homological degree and $t_{i}^{B}(B/I)$ be as defined in Section 1. Then for $a_{max}^{i}=\text{max}\{t_{i+1}^{B}(B/I)+d,(i-1)d\}$, from Remark $(1)$, $\mathcal{R}(I)_{\Delta}$ is Koszul if $\dfrac{a_{max}^{i}}{c}\leq i+1$ for all $i\geq 1$.
Case 1: $a_{max}^{i}=(i-1)d$. Then for $c\geq d$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
Case 2: $a_{max}^{i}=t_{i+1}^{B}(B/I)-d$. Then for $c\geq\dfrac{t_{i+1}^{B}(B/I)-d}{i+1}$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
Let $\gamma=\text{max}\{\dfrac{t_{i+1}^{B}(B/I)-d}{i+1}|\,1\leq i\leq\operatorname{proj\,dim}_{B}(B/I)\}$. Then for $c\geq\max\{\gamma,d\}$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
As a consequence of the discussion above, following can be observed regarding the Koszulness of the diagonals of $\mathcal{R}(I)$ for the class of ideals discussed in Section 3.
Proposition 4.1.
Let $I\subseteq B$ be a graded almost complete intersection ideal of $B$ and $S$ be a standard bigraded polynomial $K$-algebra.
  1. (1)
    Consider the assumptions as in Setup 2 with $d_{1}=d_{2}=d_{3}=d$, $d>1$ and $I$ being linearly presented. Then $\mathcal{R}(I)_{\Delta}$ is Koszul for $c\geq\frac{d}{3}$ and $e>0$.
  2. (2)
    Consider the assumptions as in Subsection 3.3$(a)$ and $(b)$. Then for $c\geq\big{(}\frac{n-1}{n}\big{)}d$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
Remark 4.2.
If $I$ is an almost complete intersection ideal of grade $2$ and perfect generated in degree $1$, then $\mathcal{R}(I)_{\Delta}$ is Koszul for all $\Delta$.
The previous proposition thus gives bounds on the diagonals of Rees algebra associated to almost complete intersections for which they are Koszul, similar to the ones given in [9, 6, 23] for Rees algebras corresponding to complete intersections.
4.2. Cohen-Macaulayness of $\mathcal{R}(I)_{\Delta}$
The following result gives the diagonals for which the diagonal subalgebras of the Rees algberas corresponding to ideals discussed in Section 3 would be Cohen-Macaulay.
Proposition 4.3.
Let $I$ be a graded ideal of $B$ satisfying the conditions in any of the subsections 3.1, 3.2 or 3.3. Then $\mathcal{R}(I)$ is Cohen-Macaulay and $\mathcal{R}(I)_{\Delta}$ is Cohen-Macaulay for all $e>0$ and $c>\max\{\min\{(e-1)d+u-m,e(u-m)\},\min\{(e-1)d+u-d_{1},e(u-d_{1})\},ed\}$.
Proof.
We have $\dim S=m+n+1=\operatorname{depth}S$ and $\dim\mathcal{R}(I)=m+1$. Since $\operatorname{proj\,dim}_{B}(B/I)=n$ in all the cases, from Remark 3.4 and Lemma 1.2, one obtains $\operatorname{depth}\mathcal{R}(I)\geq m+1$. Hence $\mathcal{R}(I)$ is Cohen-Macaulay. Then, Cohen-Macaulayness of $\mathcal{R}(I)_{\Delta}$ for the mentioned diagonals can be seen as a consequence of Remark 1.4 $(2)$. ∎
Acknowledgement. The Core Research Grant (CRG/2023/007668) from the Science and Engineering Research Board, India, partially supports the first author. INSPIRE Fellowship, DST, India, financially supports the second author.

### v2

@@
 An almost complete intersection ideal can be seen as a $d$-sequence ideal with the minimal number of generators being one more than its height.
-In this paper, we give exact formulas for the regularity of powers of graded almost complete intersection ideals.
-We also present the minimal bigraded free resolutions of Rees algebras associated with linear type ideals having one generator more than their grade and study the properties of Cohen-Macaualyness, Koszulness associated to their diagonals.
+In this paper, we give exact formulas for the regularity of powers of graded almost complete intersection ideals satisfying certain conditions.
+We also present the minimal bigraded free resolutions of Rees algebras associated with linear type ideals having one generator more than their grade and study the properties of Cohen-Macaulayness, Koszulness associated to their diagonals.
 Key words and phrases:
@@
 Primary 13D02, 13J30,13C05; Secondary 13H10
+Date: Submitted 16/03/2024, Revised: 21/10/2024
 Introduction
@@
 In [25], Rees algebras of certain Pfaffians ideals generated by $d$-sequence which come corresponding to alternating matrices are discussed.
-There are other special cases for which the correponding Rees algebras are studied [11, 33, 34, 28, 14, 26].
+There are other special cases for which the corresponding Rees algebras are studied [12, 33, 34, 28, 15, 26].
 Another class of ideals of linear type are almost complete intersections.
 An ideal $I$ in a standard graded $K$-algebra $A$ is said to be an almost complete intersection if the minimal number of generators of $I$ is one more than the height of $I$ along with the property that for all minimal primes $\mathfrak{p}$ of $I$ in $A$, $I_{\mathfrak{p}}$ is a complete intersection.
-It is proved in [21] that these ideals are contained in the class of $d$-sequence ideals.
+It is proved in [22] that homogeneous almost complete intersection ideals are contained in the class of $d$-sequence ideals (Proposition 1.5).
 Further some study has been done on their regularity [21, 31], combinatorial characterization of graphs which correspond to these ideals ([22]) and their corresponding Rees algebras [14, 22].
@@
 Finding these resolutions corresponding to arbitrary ideals in general, is a difficult problem.
-Applying stronger conditions on the ideal helps in a better understanding of the homological invariants associated to them.
-In Section 3 (Proposition 3.1) we give the minimal bigraded free resolution of the Rees algebra corresponding to an ideal satisfying the following properties: $(1)$ ideal is of linear type, $(2)$ minimal number of generators of the ideal is one more than its grade.
+Applying stronger conditions on the ideal helps in a better understanding of the homological invariants associated to it.
+In Section 3 (Theorem 3.1) we give the minimal bigraded free resolution of the Rees algebra corresponding to an ideal satisfying the following properties: $(1)$ ideal is of linear type, $(2)$ minimal number of generators of the ideal is one more than its grade.
 This is a generalization of a result in [5] extending it to the case of non-equigenerated ideals.
@@
 In particular, it shows that in order to obtain the minimal free resolution of the Rees algebra of almost complete intersection ideals, it suffices to have information about the resolution of $I$.
-A significant algebraic invariant which measures the complexity of a module is Castelnuovo-Mumford regularity.
+A significant algebraic invariant which measures the complexity of a module is the Castelnuovo-Mumford regularity.
 It is proved in [11, 23] that for every homogeneous ideal $I$, the regularity of $I^{s}$ is of the form $as+b$ for $s\gg 0$ and non-negative constants $a,b$.
@@
 In Section 4, we study the ring-theoretic properties like Cohen-Macaulayness, Koszulness of the diagonals of Rees algebras associated to certain specific classes of ideals of linear type.
-For a complete intersection ideal $I$ in a polynomial ring minimally generated by $r$ elements of degree $d$, it is proved in [9, Corollary 6.10] that for $c\geq\frac{d(r-1)}{r}$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul and an improved bound of $c\geq\frac{d}{2}$ is given for the case of $r=2$, $d=2$ in [6, Corollary 3.3] and for the general case of $r=2$ and any $d$ in [23, Theorem 3.1]).
+For a complete intersection ideal $I$ in a polynomial ring minimally generated by $r$ elements of degree $d$, it is proved in [10, Corollary 6.10] that for $c\geq\frac{d(r-1)}{r}$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul and an improved bound of $c\geq\frac{d}{2}$ is given for the case of $r=2$, $d=2$ in [7, Corollary 3.3] and for the general case of $r=2$ and any $d$ in [24, Theorem 3.1]).
 In Proposition 4.1$(1)$, we prove a similar result for an almost complete intersection ideal $I$ minimally generated by $3$ elements in degree $d$ in a graded polynomial ring, which is, $\mathcal{R}(I)_{\Delta}$ is Koszul for $c\geq\frac{d}{3}$ and $e>0$.
 1. Preliminaries
-Following are some notations used throughout the article.
+The following are some notations used throughout the article.
   1. a)
@@
 $\mathbb{F}:\cdots\stackrel{{\scriptstyle d_{\mathnormal{i+1}}}}{{\longrightarrow}}\bigoplus_{{\mathnormal{j}}}A(-\mathnormal{j})^{{{\beta_{\mathnormal{i},\mathnormal{j}}}}}\stackrel{{\scriptstyle d_{\mathnormal{i}}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle d_{\mathnormal{2}}}}{{\longrightarrow}}\bigoplus_{{\mathnormal{j}}}A(-\mathnormal{j})^{{{\beta_{\mathnormal{1},\mathnormal{j}}}}}\stackrel{{\scriptstyle d_{\mathnormal{1}}}}{{\longrightarrow}}\bigoplus_{{\mathnormal{j}}}A(-\mathnormal{j})^{{{\beta_{\mathnormal{0},\mathnormal{j}}}}}\longrightarrow 0.$
-Set $t_{\mathnormal{i}}^{A}(M):=\sup\{\mathnormal{j}:\beta_{\mathnormal{i},\mathnormal{j}}^{A}(M)\neq 0\}$ where $\beta_{i,j}$ is the $(i,j)^{\text{th}}$ Betti number of $M$ and by convention, $t_{\mathnormal{i}}^{A}(M)=-\infty$ if $F_{\mathnormal{i}}=0.$ Castelnuovo-Mumford regularity is then defined as follows:
+Set $t_{\mathnormal{i}}^{A}(M):=\sup\{\mathnormal{j}:\beta_{\mathnormal{i},\mathnormal{j}}^{A}(M)\neq 0\}$ where $\beta_{i,j}$ is the $(i,j)^{\text{th}}$ Betti number of $M$ and by convention, $t_{\mathnormal{i}}^{A}(M)=-\infty$ if $F_{\mathnormal{i}}=0.$ The Castelnuovo-Mumford regularity is then defined as follows:
 $\operatorname{reg}_{A}(M)=sup\{j-i:\beta_{\mathnormal{i},\mathnormal{j}}^{A}(M)\neq 0;\,i,j\in\mathbb{N}\}=sup\{t_{i}^{A}(M)-i:i\in\mathbb{N}\}.$
-Castelnouvo-Mumford regularity measures the complexity of the minimal free resolution of a finitely generated $A$-module $M$. We use the notation $\operatorname{reg}(M)$ when the ring $A$ is evident.
+The Castelnuovo-Mumford regularity measures the complexity of the minimal free resolution of a finitely generated $A$-module $M$.
+We use the notation $\operatorname{reg}(M)$ when the ring $A$ is evident.
 Similarly, considering the bigraded minimal free resolution of a finitely generated bigraded module $N$ over a bigraded ring $R$, one obtains biregularities. They are defined as follows,
@@
 Lemma 1.1 (Regularity Lemma).
-[17, Lemma 3.1] Let $0\rightarrow M\rightarrow N\rightarrow P\rightarrow 0$ be a short exact sequence of finitely generated graded $A$-modules. Then, the following holds.
+[18, Lemma 3.1] Let $0\rightarrow M\rightarrow N\rightarrow P\rightarrow 0$ be a short exact sequence of finitely generated graded $A$-modules. Then, the following holds.
   1. (1)
@@
 Note 1.3.
-Following are some properties of $d$-sequences and regular sequences which will be used in this article.
+The following are some properties of $d$-sequences and regular sequences that will be used in this article.
   1. (1)
-    If $\{f_{1},\ldots,f_{n}\}$ forms a $d$-sequence in a ring $A$, then
-    1. (a)
-      $(\langle f_{1},\ldots,f_{n-1}\rangle:f_{n}^{s})=(\langle f_{1},\ldots,f_{n-1}\rangle:f_{n})$ for all $s\geq 1$.
-    2. (b)
-      For $s\geq 1$, ${(\langle f_{1},\ldots,f_{i-1}\rangle+I^{s}:f_{i}})=(\langle f_{1},\ldots,f_{i-1}\rangle:f_{i})+I^{s-1}$ for all $1\leq i\leq n$ ([31, Observation 2.4]).
+    Let $\{f_{1},\ldots,f_{n}\}$ be a $d$-sequence in a ring $A$. Then, $(\langle f_{1},\ldots,f_{n-1}\rangle:f_{n}^{s})=(\langle f_{1},\ldots,f_{n-1}\rangle:f_{n})$ for all $s\geq 1$.
+  2. (2)
+    Let $\{f_{1},\ldots,f_{n}\}$ be a $d$-sequence in a ring $A$, and let $I=\langle f_{1},\ldots,f_{n}\rangle$. Then, for $s\geq 1$,
+    ${(\langle f_{1},\ldots,f_{i-1}\rangle+I^{s}:f_{i})}=(\langle f_{1},\ldots,f_{i-1}\rangle:f_{i})+I^{s-1}$
+    for all $1\leq i\leq n$ ([31, Observation 2.4]).
   3. (3)
@@
 Remark 1.4.
-Following are some results related to the Koszulness and Cohen-Macaulayness of the diagonals of bigraded algebras which will be used in the paper.
+The following are some results related to the Koszulness and Cohen-Macaulayness of the diagonals of bigraded algebras which will be used in the paper.
   1. (1)
-    For $\Delta=(c,e)$ and a bigraded $K$-algebra $R$ with the $i^{th}$ Betti number corresponding to the shift $(a,b)$ being denoted by $\beta_{i,a,b}$, $R_{\Delta}$ is Koszul if $\max\{\dfrac{a}{c},\dfrac{b}{e}:\beta_{i,a,b}\neq 0\}\leq i+1$ for every $i$ ([9, Theorem 6.2]).
+    For $\Delta=(c,e)$ and a bigraded $K$-algebra $R$ with the $i^{th}$ Betti number corresponding to the shift $(a,b)$ being denoted by $\beta_{i,a,b}$, $R_{\Delta}$ is Koszul if $\max\{\dfrac{a}{c},\dfrac{b}{e}:\beta_{i,a,b}\neq 0\}\leq i+1$ for every $i$ ([10, Theorem 6.2]).
   2. (2)
@@
 If $\mathcal{R}(I)$ is Cohen-Macaulay, then for $\alpha=\text{min}\{(e-1)d+u-m,e(u-m)\}$ and $\beta=\text{min}\{(e-1)d+u-d_{1},e(u-d_{1})\}$, $\mathcal{R}(I)_{\Delta}$ is Cohen-Macaulay if $c>\text{max}\{\alpha,\beta,de\}$ [27, Theorem 4.5].
+Proposition 1.5.
+[22, Proposition 4.10] Let $I$ be a homogeneous almost complete intersection in a polynomial ring over an infinite field, with the height of $I$ denoted by $n$.
+Then, $I$ is generated by a homogeneous $d$-sequence $\{f_{1},\ldots,f_{n+1}\}$ such that $\{f_{1},\ldots,f_{n}\}$ is a regular sequence.
 2. Regularity of powers of Almost Complete Intersections
 In [4], the authors give the regularity of powers of an equigenerated complete intersection $I$ generated by a regular sequence of length $n$ and in degree $d$ as,
-$\operatorname{reg}(I^{s})=ds+(d-1)(n-1).$ (1)
+$\operatorname{reg}(I^{s})=ds+(d-1)(n-1).$
 Further, in [31], Shen and Zhu give exact formulas for the regularity of powers of equigenerated almost complete intersections.
 In this section, we extend their result to give the regularity of powers of almost complete intersections in the non-equigenerated case as well.
-Our result coincides with the result in [31] when the ideal is equigenerated.
-Note that an upper bound for the same is given in [20].
+Our result coincides with [31, Theorem 3.1] when the ideal is equigenerated.
+Note that an upper bound for the same is discussed in [21, Corollary 2.11].
 Setup 1.
@@
 Lemma 2.1.
-Consider the notation as assumptions as in Setup 1. Then,
+Consider the notation and assumptions as in Setup 1. Then, for $i=0,\ldots,n-1$,
 $\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{2}}\right)=\sum_{l=1}^{n}d_{l}-n.$
@@
 $0\longrightarrow\dfrac{A}{J:f_{n}}(-d_{n})\stackrel{{\scriptstyle.f_{n}}}{{\longrightarrow}}\dfrac{A}{J}\longrightarrow\dfrac{A}{I}\longrightarrow 0$
-Here, $\operatorname{reg}\left(\dfrac{A}{J}\right)=\sum_{l=1}^{n-1}d_{l}-(n-1)>\operatorname{reg}\left(\dfrac{A}{I}\right)$. Hence,
+Since $J$ is generated by a regular sequence $\{f_{1},\ldots,f_{n-1}\}$ of degrees $d_{1},\ldots,d_{n-1}$, respectively, the Koszul complex corresponding to the sequence $\{f_{1},\ldots,f_{n-1}\}$ resolves the ideal $J$, and it has the following form:
+$0\longrightarrow A(-(d_{1}+d_{2}+\ldots+d_{n-1}))\stackrel{{\scriptstyle\phi_{n-1}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\phi_{k+1}}}{{\longrightarrow}}\bigoplus_{{1\leq i_{1}\leq i_{2}\leq\cdots\leq i_{k}\leq n-1}}A(-(d_{i_{1}}+d_{i_{2}}+\ldots+d_{i_{k}}))\stackrel{{\scriptstyle\phi_{k}}}{{\longrightarrow}}\ldots\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\bigoplus_{j=1}^{n-1}A(-d_{j})\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}A\longrightarrow 0,$ (1)
+where $\phi_{i}^{\prime}s$ are the differentials in the Koszul complex.
+[…]
+$\operatorname{reg}\left(\dfrac{A}{J}\right)=\sum_{l=1}^{n-1}d_{l}-(n-1),$
+which is greater than $\operatorname{reg}\left(\dfrac{A}{I}\right)$ by assumption. Hence,
 $\operatorname{reg}\left(\dfrac{A}{J:f_{n}}\right)=\sum_{l=1}^{n-1}d_{l}-(n-1)-d_{n}.$ (2)
@@
 $0\longrightarrow\dfrac{A}{(J:f_{n}^{2})}(-2d_{n})\stackrel{{\scriptstyle.f_{n}^{2}}}{{\longrightarrow}}\dfrac{A}{J}\longrightarrow\dfrac{A}{J+\langle f_{n}^{2}\rangle}\longrightarrow 0$
-As a consequence of Note 1.3, $\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}^{2}\rangle}\right)=\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}\rangle}\right)=\sum_{l=1}^{n-1}d_{l}-n+1-d_{n}.$ Then $\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}^{2}\rangle}(-2d_{n})\right)=\operatorname{reg}\left(\dfrac{A}{J}\right)+d_{n}$.
+As a consequence of Note 1.3(1), we have $\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}^{2}\rangle}\right)=\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}\rangle}\right)=\sum_{l=1}^{n-1}d_{l}-n+1-d_{n}.$ Then,
+$\operatorname{reg}\left(\dfrac{A}{J:\langle f_{n}^{2}\rangle}(-2d_{n})\right)=\operatorname{reg}\left(\dfrac{A}{J}\right)+d_{n}$.
 Therefore, from Lemma 1.1(3), $\operatorname{reg}\left(\dfrac{A}{J+\langle f_{n}^{2}\rangle}\right)=\sum_{l=1}^{n}d_{l}-n$.
@@
 $0\longrightarrow\dfrac{A}{(\langle f_{1},\ldots,f_{i-1}\rangle+I^{2}):f_{i}}(-d_{i})\stackrel{{\scriptstyle.f_{i}}}{{\longrightarrow}}\dfrac{A}{\langle f_{1},\ldots,f_{i-1}\rangle+I^{2}}\longrightarrow\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{2}}\longrightarrow 0$
-Since $\{f_{1},\ldots,f_{n}\}$ forms a $d$-sequence with $\{f_{1},\ldots,f_{n-1}\}$ forming a regular sequence, from Note 1.3,
-$\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i-1}\rangle+I^{2}):f_{i}}\right)(-d_{i})=\operatorname{reg}\left(\dfrac{A}{I}\right)+d_{i}<\sum_{l=1}^{n-1}d_{i}-(n-1)+d_{i}\leq\sum_{l=1}^{n}d_{i}-(n-1)\leq$
-$\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{2}}\right)+1$. Hence from Lemma $(1)$,
+Since $\{f_{1},\ldots,f_{n}\}$ forms a $d$-sequence with $\{f_{1},\ldots,f_{n-1}\}$ forming a regular sequence, from Note 1.3(2), we have $\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i-1}\rangle+I^{2}):f_{i}}\right)(-d_{i})=\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i-1}\rangle:f_{i})+I}\right)(-d_{i}).$ From Note 1.3(3), this is equal to $\operatorname{reg}\left(\dfrac{A}{I}\right)+d_{i}<\sum_{l=1}^{n-1}d_{l}-(n-1)+d_{i}\leq\sum_{l=1}^{n}d_{l}-(n-1)\leq\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{2}}\right)+1$.
+Hence, from Lemma 1.1(1),
 $\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i-1}\rangle+I^{2}}\right)=\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{2}}\right)=\sum_{l=1}^{n}d_{l}-n.$
@@
 $0\longrightarrow\dfrac{A}{J:f_{n}^{s+1}}(-sd_{n})\stackrel{{\scriptstyle.f_{n}^{s}}}{{\longrightarrow}}\dfrac{A}{J:f_{n}}\longrightarrow\dfrac{A}{(J:f_{n})+\langle f_{n}^{s}\rangle}\longrightarrow 0$
-From Note 1.3, one obtains $\operatorname{reg}\left(\dfrac{A}{J:f_{n}^{s+1}}\right)(-sd_{n})=\operatorname{reg}\left(\dfrac{A}{J:f_{n}}\right)+sd_{n}$. Then from Lemma 1.1$(3)$,
+From Note 1.3(1), one obtains $\operatorname{reg}\left(\dfrac{A}{J:f_{n}^{s+1}}\right)(-sd_{n})=\operatorname{reg}\left(\dfrac{A}{J:f_{n}}\right)+sd_{n}$. Then from Lemma 1.1(3),
 $\operatorname{reg}\left(\dfrac{A}{(J:f_{n})+\langle f_{n}^{s}\rangle}\right)=\operatorname{reg}\left(\dfrac{A}{J:f_{n}}\right)+sd_{n}-1.$ (3)
 Now substituting the value of $\operatorname{reg}\left(\dfrac{A}{J:f_{n}}\right)$ from Equation (2) in Equation (3), we get the required result.
-As a consequence of the above results, one can obtain the regularity of powers of almost complete intersections.
+From Proposition 1.5, it is known that homogeneous almost complete intersection ideals in a polynomial ring over an infinite field have a generating set of $d$-sequence such that removing the last generator results in a regular sequence.
+Hence, as a consequence of Lemmas 2.1 and 2.2, one can determine the regularity of powers of almost complete intersections that satisfy certain conditions.
 Theorem 2.3.
-Consider the notation and assumptions as in Setup 1. Then for all $s\geq 2$ and $i=0,\ldots,n-1$, the following is true,
+Let $A$ be a standard graded polynomial ring over an infinite field $K$ and $I$ be a homogeneous almost complete intersection ideal of $A$ generated by a $d$-sequence $\{f_{1},\ldots,f_{n}\}$ in degrees $d_{1},\ldots,d_{n}$, where $\{f_{1},\ldots,f_{n-1}\}$ is a regular sequence and $J=\langle f_{1},\ldots,f_{n-1}\rangle$.
+Assume that $\operatorname{reg}(A/I)<\sum_{l=1}^{n-1}d_{l}-n+1$ and that $d_{n}=\max\{d_{i}\mid i=1,\ldots,n\}$ Then for all $s\geq 2$ and $i=0,\ldots,n-1$, the following is true,
 $\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{s}}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}.$
@@
 Then clearly, $\langle f_{1},\ldots,f_{(n-1)+1}\rangle+I^{s}=I$.
-From Note 1.3 and Lemma 2.2,
-$\operatorname{reg}\left(\dfrac{A}{(J+I^{s}):f_{n}}(-d_{n})\right)=\operatorname{reg}\left(\dfrac{A}{(J:f_{n})+I^{s-1}}\right)+d_{n}=\sum_{l=1}^{n}d_{l}-n+(s-3)d_{n}+d_{n}=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}\geq\sum_{l=1}^{n-1}d_{l}-(n-1)+1>\operatorname{reg}\left(\dfrac{A}{I}\right)+1$.
-Hence $\operatorname{reg}\left(\dfrac{A}{J+I^{s}}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}$ by Lemma $(1)$.
-Now assume $i\leq n-2$ and that the result holds for $i+1$.
-In the short exact sequence (4), by induction hypothesis on $i$, $\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i+1}\rangle+I^{s}}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}$.
-Since $\{f_{1},\ldots,f_{j}\}$ forms a $d$-sequence for $1\leq j\leq n$ and a regular sequence for $11\leq j\leq n-$ 1, from Note 1.3, $\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i}\rangle+I^{s}):f_{i+1}}\right)=\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{s-1}}\right)$.
-This by induction hypothesis on $s$ is equal to $\sum_{l=1}^{n}d_{l}-n+(s-3)d_{n}$.
-Now, the final result is a consequence of the inequality
-$\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i}\rangle+I^{s}):f_{i+1}}\right)+d_{i+1}<\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i+1}\rangle+I^{s}}\right)+1$
-and Lemma $(1)$. ∎
+From Note 1.3(2) and Lemma 2.2,
+$\operatorname{reg}\left(\dfrac{A}{(J+I^{s}):f_{n}}(-d_{n})\right)=\operatorname{reg}\left(\dfrac{A}{(J:f_{n})+I^{s-1}}\right)+d_{n}=\operatorname{reg}\left(\dfrac{A}{(J:f_{n})+\langle f_{n}^{s-1}\rangle}\right)+d_{n}=\sum_{l=1}^{n}d_{l}-n+(s-3)d_{n}+d_{n}=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}\geq\sum_{l=1}^{n-1}d_{l}-(n-1)+1>\operatorname{reg}\left(\dfrac{A}{I}\right)+1$.
+Hence $\operatorname{reg}\left(\dfrac{A}{J+I^{s}}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}$ by Lemma 1.1(1).
+Now assume $i\leq n-2$ and that the result holds for $i+1$. In the short exact sequence (4), by induction hypothesis on $i$,
+$\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i+1}\rangle+I^{s}}\right)=\sum_{l=1}^{n}d_{l}-n+(s-2)d_{n}.$
+Since $\{f_{1},\ldots,f_{j}\}$ forms a $d$-sequence for $1\leq j\leq n$ and a regular sequence for $11\leq j\leq n-$ 1, from Note 1.3(2) and Note 1.3(3), $\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i}\rangle+I^{s}):f_{i+1}}\right)=\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i}\rangle+I^{s-1}}\right)$.
+This, by induction hypothesis on $s$, is equal to $\sum_{l=1}^{n}d_{l}-n+(s-3)d_{n}$.
+Now, the final result is a consequence of the inequality,
+$\operatorname{reg}\left(\dfrac{A}{(\langle f_{1},\ldots,f_{i}\rangle+I^{s}):f_{i+1}}\right)+d_{i+1}\leq\sum_{l=1}^{n}d_{l}-n+(s-3)d_{n}+d_{n}<\operatorname{reg}\left(\dfrac{A}{\langle f_{1},\ldots,f_{i+1}\rangle+I^{s}}\right)+1$
+and Lemma 1.1(1). ∎
 Note 2.4.
@@
 But $\operatorname{reg}\left(\dfrac{B}{I^{3}}\right)=8$ which is not equal to $(2+2+2+3)-4+2=7$.
-As a consequence of the above results, the following can be observed regarding the regularity of powers of almost complete intersections of grade $3$.
+Note 2.5.
+It is important to note that the formula for the regularity of powers in Theorem 2.3 holds true for a general setup of a homogeneous ideal $I$ in a standard graded $K$-algebra $A$, where the field is not necessarily infinite, provided that the following properties are satisfied:
+  1. (1)
+    $I$ has a generating set of $d$-sequence $\{f_{1},\ldots,f_{n}\}$ such that $\{f_{1},\ldots,f_{n-1}\}$ is a regular sequence.
+  2. (2)
+    $\operatorname{reg}(A/I)<\sum_{l=1}^{n-1}d_{l}-n+1$ and $d_{n}=\max\{d_{i}\mid i=1,\ldots,n\}$.
+As a consequence of Theorem 2.3 and Note 2.5, the following can be observed regarding the regularity of powers of almost complete intersections of grade $3$.
 To understand the set of generators of these ideals, we look at the notion of Pfaffians.
 Given a skew-symmetric matrix $X$, the Pfaffian of $X$ is defined as the square root of the determinant of $X$.
-For $l\in\mathbb{N}$, let $\operatorname{Pf}_{\bar{l}}(X)$ denotes the Pfaffian of $X$ obtained by removing the $l^{th}$ row and the corresponding column of $X$ which can be seen as an element in $B=K[X]$.
-In [7], the authors show that almost complete intersections of grade $3$ arise as the ideal generated by certain Pfaffians of skew-symmetric matrices of some particular form (see Setup 3).
-Following are some properties of these ideals that can be observed.
+For $l\in\mathbb{N}$, let $\operatorname{Pf}_{\bar{l}}(X)$ denote the Pfaffian of $X$ obtained by removing the $l^{th}$ row and the corresponding column of $X$, which can be seen as an element in $B=K[X]$.
+Let $X$ be skew-symmetric matrix of order $t$ with the following form:
+$X=\begin{bmatrix}0&0&0&x_{14}&x_{15}&\cdots&x_{1t}\\ 0&0&0&x_{24}&x_{25}&\cdots&x_{2t}\\ 0&0&0&x_{34}&x_{35}&\cdots&x_{3t}\\ -x_{14}&-x_{24}&-x_{34}&0&x_{45}&\cdots&x_{4t}\\ -x_{15}&-x_{25}&-x_{35}&-x_{45}&0&\cdots&x_{5t}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\cdots&\vdots\\ -x_{1t}&-x_{2t}&-x_{3t}&-x_{4t}&-x_{5t}&\cdots&0\\ \end{bmatrix}$ (5)
+Then, from the results and discussions in [6] and [8], it can be concluded that an almost complete intersection ideal $I$ in $K[X]$ of grade $3$ and type at least $2$ arises as an ideal generated by Pfaffians of the following form,
   1. (1)
-    $B/I$ is Cohen-Macaulay. Since $I$ is perfect in $B$, the observation follows from [12, Section 19].
+    If $t\geq 5$ is an odd integer, then
+    $I=\langle\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X),\operatorname{Pf}_{\overline{123}}(X)\rangle.$ (6)
   2. (2)
-    For $t\geq 5$ and $I$ an ideal of type $t-3$,
-    1. (a)
-      $\text{reg}_{B}(B/I)=t-4$.
-    2. (b)
-      $(B/I)^{(c)}$ is Koszul for $c\geq\dfrac{t}{3}$.
+    If $t\geq 6$ is an even integer, then
+    $I=\langle\operatorname{Pf}(X),\operatorname{Pf}_{\overline{12}}(X),\operatorname{Pf}_{\overline{13}}(X),\operatorname{Pf}_{\overline{23}}(X)\rangle.$ (7)
+Setup 2.
+Let $B=K[X]$ denote a standard graded polynomial ring with $X$ being a skew-symmetric matrix of the form (5).
+For $t\in\mathbb{N}$ with $t\geq 5$, let $I$ be an almost complete intersection ideal of $B$ of grade $3$ and type $t-3$ (Equations (6) and (7)).
+Remark 2.6.
+Note that from [8, Theorems 1.2 and 1.3], one can obtain the minimal free resolutions of ideals of the form in Setup 2. As a consequence, the following properties can be observed.
+  1. (1)
+    $B/I$ is Cohen-Macaulay. Since $I$ is perfect in $B$, the observation follows from [13, Section 19].
+  2. (2)
+    For $t\geq 5$ and $I$ an ideal of type $t-3$, $\text{reg}(B/I)=t-4$.
+  3. (3)
+    For $t\geq 5$ and $I$ an ideal of type $t-3$, $(B/I)^{(c)}$ is Koszul for $c\geq\frac{t}{3}$.
 Lemma 2.7.
-Assume the notation and hypotheses as in Setup 3. Then,
+Let $X$ be a skew-symmetric matrix of order $t$ of the form given in Equation (5). Then,
   1. (1)
-    For an odd integer $t\geq 5$, $\{\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X)\}$ is a $d$-sequence generating set of an almost complete intersection ideal of grade $3$ and type $t-3$.
+    For an odd integer $t\geq 5$, $\{\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X)\}$ is a $d$-sequence such that $\{\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X)\}$ forms a regular sequence in $K[X]$.
   2. (2)
-    For an even integer $t\geq 6$, $\{\operatorname{Pf}_{\overline{12}}(X),\operatorname{Pf}_{\overline{13}}(X),\operatorname{Pf}_{\overline{23}}(X),\operatorname{Pf}(X)\}$ is a $d$-sequence generating set of an almost complete intersection ideal of grade $3$ and type $t-3$.
+    For an even integer $t\geq 6$, $\{\operatorname{Pf}_{\overline{12}}(X),\operatorname{Pf}_{\overline{13}}(X),\operatorname{Pf}_{\overline{23}}(X),\operatorname{Pf}(X)\}$ is a $d$-sequence such that $\{\operatorname{Pf}_{\overline{12}}(X),\operatorname{Pf}_{\overline{13}}(X),\operatorname{Pf}_{\overline{23}}(X)\}$ forms a regular sequence in $K[X]$.
 Proof.
-We have $\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X)\}$ forms a regular sequence from [7, C.2 Lemma].
-Further, since $J$ is unmixed, from [13, Proposition 1.3], $\{\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X)\}$ forms a $d$-sequence.
-Similarly $(2)$ can be seen as a consequence of [7, C.6 Lemma] and [13, Proposition 1.3]. ∎
+We have $\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X)\}$ forms a regular sequence from [8, C.2 Lemma].
+Further, since $J$ is unmixed, from [14, Proposition 1.3], $\{\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X)\}$ forms a $d$-sequence.
+Similarly $(2)$ can be seen as a consequence of [8, C.6 Lemma] and [14, Proposition 1.3]. ∎
 Corollary 2.8.
-Assume the notation and hypotheses as in Setup 3. Then for $m\geq 2$,
+Assume the notation and hypotheses as in Setup 2. Then for $m\geq 2$,
   1. (1)
-    If $t\geq 5$ is an odd integer such that , then for $i=1,2,3$,
-    $\operatorname{reg}\left(\dfrac{B}{\langle f_{1},\ldots,f_{i}\rangle+I^{m}}\right)=\left(\frac{t-1}{2}\right)(m+2)-5.$
+    If $t\geq 5$ is an odd integer, then for $i=1,2,3$,
+    $\operatorname{reg}\left(\dfrac{B}{I^{m}}\right)=\left(\frac{t-1}{2}\right)(m+2)-5.$
   2. (2)
     If $t\geq 6$ is an even integer, then for $i=1,2,3$,
-    $\operatorname{reg}\left(\dfrac{B}{\langle f_{1},\ldots,f_{i}\rangle+I^{m}}\right)=\left(\frac{t}{2}\right)(m+2)-7.$
+    $\operatorname{reg}\left(\dfrac{B}{I^{m}}\right)=\left(\frac{t}{2}\right)(m+2)-7.$
 Proof.
-The result can be seen as a consequence of Theorem 2.3 and Lemma 2.5
+We have $I$ to be an almost complete intersection ideal of $B=K[X]$, a standard graded polynomial ring with $X$ being a skew-symmetric matrix of the form (5), of grade $3$ and type $t-3$ for $t\in\mathbb{N}$ and $t\geq 5$.
+  1. (1)
+    If $t$ is an odd integer, then from Equation (6), $I$ is generated by the Pfaffians $\operatorname{Pf}_{\overline{123}}(X),\operatorname{Pf}_{\bar{1}}(X),\operatorname{Pf}_{\bar{2}}(X),\operatorname{Pf}_{\bar{3}}(X)$ which are of degrees $\frac{t-3}{2},\frac{t-1}{2},\frac{t-1}{2},\frac{t-1}{2}$, respectively.
+From Remark 2.6(2), we have $\text{reg}(B/I)=t-4<\left(\frac{t-3}{2}\right)+\left(\frac{t-1}{2}\right)+\left(\frac{t-1}{2}\right)-4+1$.
+Thus, the result can be seen as a consequence of Lemma 2.7, Note 2.5 and the simplification of the formula derived from Theorem 2.3.
+  2. (2)
+    If $t$ is an even integer, then from Equation (7), $I$ is generated by the Pfaffians $\operatorname{Pf}_{\overline{12}}(X),\operatorname{Pf}_{\overline{13}}(X),\operatorname{Pf}_{\overline{23}}(X),\operatorname{Pf}(X)$ which are of degrees $\frac{t-2}{2},\frac{t-2}{2},\frac{t-2}{2},\frac{t}{2}$, respectively.
+From Remark 2.6(2), since $\text{reg}(B/I)=t-4<\left(\frac{t-2}{2}\right)+\left(\frac{t-2}{2}\right)+\left(\frac{t-2}{2}\right)-4+1$, similar to the previous case, the result follows from Lemma 2.7, Note 2.5 and Theorem 2.3.
 3. Rees algebra of Almost Complete Intersections
 This section discusses the resolutions of the Rees algebra of some classes of ideals $I$ of linear type satisfying a relation between their minimal number of generators and their respective heights.
-The resolution is obtained as the mapping cone between the resolution of $I$ and the Buchsbaum-Rim complex, coming corresponding to the generators of $I$.
+The resolution is obtained as the mapping cone between the resolution of $I$ and the Buchsbaum-Rim complex, corresponding to the generators of $I$.
 The idea has been discussed in [5] for constructing resolutions of equigenerated ideals.
@@
 As a consequence of this, minimal bigarded free resolutions of almost complete intersections are obtained.
-The following are the building blocks for constructing the minimal bigraded free resolutions given in Proposition 3.1.
+The following are the building blocks for constructing the minimal bigraded free resolutions given in Theorem 3.1.
 Let $I$ be an ideal of a standard graded ring $A$ generated by $\{f_{1},\ldots,f_{l}\}$ with $\deg(f_{i})=d_{i}$ and grade $I=l-1$.
@@
 Then,
-    $\epsilon(e_{i})=\sum_{m\subset k,\,|m|=2}\text{sgn}(m\subset k)(\text{det}(\psi_{m}))e_{k\backslash m}$
+    $\epsilon(e_{i})=\sum_{m\subset k,\,|m|=2}\text{sgn}(m\subset k)(\text{det}(\psi_{m}))e_{k\backslash m}$ (8)
     where $\text{sgn}(m\subset k)$ is the sign of the permutation of $k$ which puts the elements of $m$ in the first $2$ positions and $\psi_{m}$ is the minor of order $2$ with the columns indexed by the elements of $m$ [5].
@@
     For $22\leq i\leq l-$ 2, the maps $\sigma_{i}$ are defined as follows,
-    $\sigma_{i}(\beta\otimes\gamma)=\sum_{s,t}[\psi^{*}(\beta_{s}^{\prime})](\gamma_{t}^{\prime})\cdot\gamma_{t}^{\prime\prime}\otimes\beta_{s}^{\prime\prime}.$
+    $\sigma_{i}(\beta\otimes\gamma)=\sum_{s,t}[\psi^{*}(\beta_{s}^{\prime})](\gamma_{t}^{\prime})\cdot\gamma_{t}^{\prime\prime}\otimes\beta_{s}^{\prime\prime}.$ (9)
 where $\beta\in(\operatorname{Sym}_{i-1}G)^{*}$ with $\beta=\sum_{s}\beta_{s}^{\prime}\otimes\beta_{s}^{\prime\prime}\in(G)^{*}\otimes(\operatorname{Sym}_{i-2}G)^{*}$ and $\gamma\in\wedge^{i+2}{F}$ with $\gamma=\sum_{s}\gamma_{t}^{\prime}\otimes\gamma_{t}^{\prime\prime}\in\wedge^{1}F\otimes\wedge^{i+1}F$.
-Next we mention a map of complexes from $\mathbb{C.}$ to $\mathbb{F.}$ mentioned in [5].
-Let $\nu$ be a map of complexes from the Koszul complex corresponding to $\phi_{1}$, denoted by (.$(\wedge^{.}A^{l},c.)$, to the Buchsbaum-Rim complex (.$(\mathbb{C.},\sigma.)$, obtained by extending the identity map at degree 00 and degree $1$ such that the diagram commutes.
-This map $\nu$ helps in defining the required map of complexes $\alpha:\mathbb{C.}\rightarrow\mathbb{F.}$ which has the following form.
-For $21\leq i\leq l-$ 2, $\alpha_{i}:\wedge^{i+2}F\otimes(\operatorname{Sym}_{i-1}G)^{*}\rightarrow F_{i+1}\otimes S$ is given as the composition of the following two maps.
-$\alpha_{i}:\wedge^{i+2}F\otimes(\operatorname{Sym}_{i-1}G)^{*}\stackrel{{\scriptstyle p\otimes\tau}}{{\longrightarrow}}\wedge^{1}F\otimes\wedge^{i+1}F\stackrel{{\scriptstyle\tiny{\begin{bmatrix}y_{1}&\cdots&y_{l}\end{bmatrix}}\otimes\nu_{i+1}}}{{\longrightarrow}}S\otimes F_{i+1}$ (5)
+For $i=1,\ldots,l$, we have $\psi^{*}(g_{1}^{*})(e_{i})=f_{i}$ and $\psi^{*}(g_{2}^{*})(e_{i})=y_{i}$.
+This clearly indicates that the way the differentials are defined means that the degrees of the generators $f_{i}$ are crucial in determining the bishifts in the Buchsbaum-Rim complex.
+Let,
+.: $\mathbb{C^{\prime}.:}\hskip 85.35826pt\;0\longrightarrow\wedge^{l}F\otimes(\text{Sym}_{l-3}G)^{*}\stackrel{{\scriptstyle\sigma_{l-2}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\sigma_{2}}}{{\longrightarrow}}\wedge^{3}{F}\otimes(\text{Sym}_{0}G)^{*}$ (10)
+denote the truncated Buchsbaum-Rim complex where $C_{i}^{\prime}=\wedge^{i+2}{F}\otimes(\text{Sym}_{i-1}G)^{*}$, and
+$\mathbb{F^{\prime}.:}\hskip 85.35826pt\ldots\stackrel{{\scriptstyle\phi_{4}^{\prime}}}{{\longrightarrow}}\bigoplus_{j\geq 3}S(-j,-1)^{\beta_{3j}}\stackrel{{\scriptstyle\phi_{3}^{\prime}}}{{\longrightarrow}}\bigoplus_{j\geq 2}S(-j,-1)^{\beta_{2j}}$ (11)
+denote the truncated complex $\mathbb{F.}\otimes S$ where the differentials $\phi_{i}^{\prime}$ denotes the base change of $\phi_{i}$ to $S$.
+Next we mention a map of complexes from $\mathbb{C^{\prime}.}$ to $\mathbb{F^{\prime}.}$ mentioned in [5].
+Let $\nu$ be a map of complexes from the Koszul complex corresponding to $\phi_{1}$, denoted by (.$(\wedge^{.}A^{l},c.)$, to $(\mathbb{F.},\phi.)$, the minimal bigraded free resolution of $I$, obtained by extending the identity map at degree 00 and degree $1$ such that the diagram commutes.
+Further, let $\nu^{\prime}$ denote the base change of $\nu$ to $S$.
+This map $\nu^{\prime}$ helps in defining the required map of complexes $\alpha:\mathbb{C^{\prime}.}\rightarrow\mathbb{F^{\prime}.}$ which has the following form.
+For $21\leq i\leq l-$ 2, $\alpha_{i}:\wedge^{i+2}F\otimes(\operatorname{Sym}_{i-1}G)^{*}\rightarrow F_{i+1}^{\prime}$ is given as the composition of the following two maps.
+$\alpha_{i}:\wedge^{i+2}F\otimes(\operatorname{Sym}_{i-1}G)^{*}\stackrel{{\scriptstyle p\otimes\tau}}{{\longrightarrow}}\wedge^{1}F\otimes\wedge^{i+1}F\stackrel{{\scriptstyle\tiny{\begin{bmatrix}y_{1}&\cdots&y_{l}\end{bmatrix}}\otimes\nu_{i+1}^{\prime}}}{{\longrightarrow}}S\otimes F_{i+1}^{\prime}$ (12)
 where $p:\wedge^{i+2}F\rightarrow\wedge^{1}F\otimes\wedge^{i+1}F$ is the co-multiplication map defined as $e_{j_{1}}\wedge e_{j_{2}}\wedge\cdots\wedge e_{j_{i+2}}\mapsto\sum_{i}e_{j_{i}}\otimes e_{j_{1}}\wedge\ldots\wedge\hat{e_{j_{i}}}\wedge\ldots\wedge e_{j_{i+2}}$ and $\tau:(\operatorname{Sym}_{i-1}G)^{*}\rightarrow S$ defined as $g_{1}^{*(i-1-j)}\otimes g_{2}^{*(j)}\mapsto\left\{\begin{array}[]{ccc}1&\text{if}&j=0\\ 0&\text{if}&j\geq 1.\\ \end{array}\right.$
 With the notations set, we move on to give the main result in this section.
-Proposition 3.1.
+Theorem 3.1.
 Let $B=K[x_{1},\ldots,x_{m}]$ be a standard graded polynomial $K$-algebra and $I=\langle f_{1},\ldots,f_{n+1}\rangle$ be a graded ideal of $B$ of linear type with deg $f_{j}=d_{j}$ for $j=1,\ldots,n+1$ and grade$(I)=n$ (ht$(I)=n$).
@@
 Then the minimal non-standard bigraded free resolution of $\mathcal{R}(I)$ has the form,
-$\cdots\stackrel{{\scriptstyle\delta_{4}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 3}S(-j,-1)^{\beta_{4j}}\\ \oplus\\ C_{2}\end{matrix}\stackrel{{\scriptstyle\delta_{3}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 2}S(-j,-1)^{\beta_{3j}}\\ \oplus\\ C_{1}\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 1}S(-j,-1)^{\beta_{2j}}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0.$ (6)
+$\cdots\stackrel{{\scriptstyle\delta_{4}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 4}S(-j,-1)^{\beta_{4j}}\\ \oplus\\ C_{2}^{\prime}\end{matrix}\stackrel{{\scriptstyle\delta_{3}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 3}S(-j,-1)^{\beta_{3j}}\\ \oplus\\ C_{1}^{\prime}\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 2}S(-j,-1)^{\beta_{2j}}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0.$ (13)
 where $\beta_{ij}$ are the $(i,j)^{th}$ Betti numbers of $B/I$ and for $i=1,\dots,n-1$,
-$C_{i}=\bigoplus_{j=0}^{i-1}\bigoplus_{1\leq j_{1}<j_{2}<\cdots<j_{i+2}\leq n+1}S(-(d_{j_{1}}+d_{j_{2}}+\cdots+d_{j_{i+2}}),-j-2)$
-are the modules which occur in the Buchsbaum-Rim complex $\mathbb{C}.$ corresponding to the matrix $\begin{bmatrix}f_{1}&\cdots&f_{n+1}\\ y_{1}&\cdots&y_{n+1}\\ \end{bmatrix}$ with differentials denoted by $\sigma_{i}$.
-Let $\alpha:\mathbb{C}.\rightarrow\mathbb{F}.$ be the map of complexes given in Equation (5).
+$C_{i}^{\prime}=\bigoplus_{j=0}^{i-1}\bigoplus_{1\leq j_{1}<j_{2}<\cdots<j_{i+2}\leq n+1}S(-(d_{j_{1}}+d_{j_{2}}+\cdots+d_{j_{i+2}}),-j-2)$
+are the modules which occur in the Buchsbaum-Rim complex $\mathbb{C^{\prime}}.$ corresponding to the matrix $\begin{bmatrix}f_{1}&\cdots&f_{n+1}\\ y_{1}&\cdots&y_{n+1}\\ \end{bmatrix}$ with differentials denoted by $\sigma_{i}$.
+Let $\alpha:\mathbb{C^{\prime}}.\rightarrow\mathbb{F^{\prime}}.$ be the map of complexes given in Equation (12).
 Then for all $i=1,\ldots,\operatorname{proj\,dim}_{S}\mathcal{R}(I)$, the differentials $\delta_{i}$ are of the following form.
-$\delta_{1}=\begin{bmatrix}y_{1}&\ldots&y_{n+1}\end{bmatrix}\circ\phi_{2},\;\;\delta_{2}=\begin{pmatrix}\alpha_{1}&\phi_{3}\end{pmatrix},\;\;\delta_{i}=\begin{pmatrix}\phi_{i+1}&\alpha_{i-1}\\ O&\sigma_{i-1}\end{pmatrix}.$
+$\delta_{1}=\begin{bmatrix}y_{1}&\ldots&y_{n+1}\end{bmatrix}\circ\phi_{2}^{\prime},\;\;\delta_{2}=\begin{pmatrix}\alpha_{1}&\phi_{3}^{\prime}\end{pmatrix},\;\;\delta_{i}=\begin{pmatrix}\phi_{i+1}^{\prime}&\alpha_{i-1}\\ O&-\sigma_{i-1}\end{pmatrix}.$
 Proof.
@@
 Here for $i=1,\ldots,n+1$, $\psi^{*}(g_{1}^{*})(e_{i})=f_{i}$ and $\psi^{*}(g_{2}^{*})(e_{i})=y_{i}$ which determines the entries in the matrices corresponding to the differentials $\sigma_{i}^{\prime}s$.
-Then the way the differentials are defined in Complex (8), one obtains
-$C_{i}=\bigoplus_{j=0}^{i-1}\bigoplus_{1\leq j_{1}<j_{2}<\cdots<j_{i+2}\leq n+1}S(-(d_{j_{1}}+d_{j_{2}}+\cdots+d_{j_{i+2}}),-j-2)$
-for $i=1,\ldots,n-1$.
+Then, the way the differentials are defined in Equation (9), one obtains
+$C_{i}^{\prime}=\bigoplus_{j=0}^{i-1}\bigoplus_{1\leq j_{1}<j_{2}<\cdots<j_{i+2}\leq n+1}S(-(d_{j_{1}}+d_{j_{2}}+\cdots+d_{j_{i+2}}),-j-2)$ (16)
+for $i=2,\ldots,n-1$.
 Consider (.$(\wedge^{.}B^{n+1},k.)$ to be the Koszul complex corresponding to $\phi_{1}$.
 Now since for $\{j_{1},j_{2}\}\subseteq[n+1]$ such that $j_{1}\neq j_{2}$, $k_{2}(e_{j_{1}}\wedge e_{j_{2}})=f_{j_{1}}e_{j_{2}}-f_{j_{2}}e_{j_{1}}\in$ Ker ϕ1=subscriptitalic-$\phi_{1}=$ Im $\phi_{2}$, there exists $\beta_{j_{1}j_{2}}\in F_{2}$ such that $\phi_{2}(\beta_{j_{1}j_{2}})=i(k_{2}(e_{j_{1}}\wedge e_{j_{2}}))$ where $i$ is the natural map from $\wedge^{0}B^{n+1}$ to $F_{1}$.
-Let $\alpha:\mathbb{C.}\rightarrow\mathbb{F.}$ be the map of complexes.
-Then $\alpha_{1}:\wedge^{3}{S^{n+1}}\otimes(\operatorname{Sym}_{0}S^{2})^{*}\rightarrow S\otimes F_{2}$ will be defined as, for $\{j_{1},j_{2},j_{3}\}\subseteq[n+1]$, $\alpha_{1}((e_{j_{1}}\wedge e_{j_{2}}\wedge e_{j_{3}})\otimes 1)=y_{j_{1}}\beta_{j_{2}j_{3}}-y_{j_{2}}\beta_{j_{1}j_{3}}+y_{j_{3}}\beta_{j_{1}j_{2}}.$
-With the details above, one obtains the required resolution (6) by considering the mapping cone of $\alpha$ (arguments similar to the ones in [5]).
+Let $\alpha:\mathbb{C^{\prime}.}\rightarrow\mathbb{F^{\prime}.}$ be the map of complexes given in Equation (12) where $\mathbb{C^{\prime}.}$ and $\mathbb{F^{\prime}.}$ are the complexes given in Equations (10) and (11) respectively.
+Then $\alpha_{1}:\wedge^{3}{S^{n+1}}\otimes(\operatorname{Sym}_{0}S^{2})^{*}\rightarrow F_{2}^{\prime}$ will be defined as, for $\{j_{1},j_{2},j_{3}\}\subseteq[n+1]$,
+$\alpha_{1}((e_{j_{1}}\wedge e_{j_{2}}\wedge e_{j_{3}})\otimes 1)=y_{j_{1}}\beta_{j_{2}j_{3}}-y_{j_{2}}\beta_{j_{1}j_{3}}+y_{j_{3}}\beta_{j_{1}j_{2}}.$
+Thus, from the way the map $\alpha_{1}$ is defined, one obtains,
+$C_{1}^{\prime}=\bigoplus_{1\leq j_{1}<j_{2}<j_{3}\leq n+1}S(-(d_{j_{1}}+d_{j_{2}}+d_{j_{3}}),-2).$ (17)
+With the details above, one obtains the required resolution (13) by considering the mapping cone of $\alpha$ and following arguments similar to those in [5].
+From the construction of the Complex (13), it is evident that the bishifts in the resolution of the Rees algebra of $I$ is derived from the information in the Complexes (14) and (15) and Equations (16) and (17).
 Further, the resolution is minimal since the entries in the matrices given by the differentials $\delta_{i}$ are from $\mathfrak{m}S$, where $\mathfrak{m}$ is the bigraded maximal ideal of $S$.
-Note that the linear shifts in the second variable is contributed by the differentials in the minimal free resolution of $B/I$ and the higher degree shifts for the same are contributed by $\alpha_{i}^{\prime}s$ and $\sigma_{i}^{\prime}s$.
+Note that the linear shifts in the second variable are contributed by the differentials in the minimal free resolution of $B/I$, while the higher degree shifts are contributed by the $\alpha_{i}^{\prime}s$ and $\sigma_{i}^{\prime}s$.
 ∎
@@
 Remark 3.3.
-Note that in Proposition 3.1 if $I$ is considered to be a graded ideal of $B$ with $\mu(I)=\text{ht}(I)+1$ and not necessarily of linear type, then the Complex (6) gives a minimal bigraded free resolution of the $\operatorname{Sym}(I)$.
+Note that in Theorem 3.1 if $I$ is considered to be a graded ideal of $B$ with $\mu(I)=\text{ht}(I)+1$ and not necessarily of linear type, then the Complex (13) gives a minimal bigraded free resolution of the $\operatorname{Sym}(I)$.
 Remark 3.4.
@@
 $\operatorname{proj\,dim}_{S}\mathcal{R}(I)=\left\{\begin{array}[]{ccc}n&\text{if}&\text{pd }(B/I)=n\\ k-1&\text{if}&\text{pd }(B/I)=k,\,k>n.\\ \end{array}\right.$
-Following observations can be made on the nature of the generators of $I$ as a consequence of the explicit minimal bigraded free resolutions of the symmetric algebra of $I$ obtained from Proposition 3.1.
-Note that for an ideal $I$ in $B$, it’s generating set $\{a_{1},\ldots,a_{n}\}$ forms a strong $s$-sequence (w.r.t an admissible term order for the monomials in $y_{i}$ with $y_{1}<y_{2}<\ldots<y_{n}$) if for $\operatorname{Sym}(I)\cong S/J$ where $S$ is a bigraded polynomial ring $S=B[y_{1},\ldots,y_{n}]$ and $J$ the defining ideal of the symmetric algebra, $in(J)=\langle L_{1}y_{1},\ldots,L_{n}y_{n}\rangle$ where $L_{i}=(a_{1},\ldots,a_{i-1}:a_{i})$ and $L_{1}\subseteq L_{2}\subseteq\ldots\subseteq L_{n}$ (cf.
+The following observations can be made on the nature of the generators of $I$ as a consequence of the explicit minimal bigraded free resolutions of the symmetric algebra of $I$ obtained from Theorem 3.1.
+Note that for an ideal $I$ in $B$, its generating set $\{a_{1},\ldots,a_{n}\}$ forms a strong $s$-sequence (w.r.t an admissible term order for the monomials in $y_{i}$ with $y_{1}<y_{2}<\ldots<y_{n}$) if for $\operatorname{Sym}(I)\cong S/J$ where $S$ is a bigraded polynomial ring $S=B[y_{1},\ldots,y_{n}]$ and $J$ the defining ideal of the symmetric algebra, $in(J)=\langle L_{1}y_{1},\ldots,L_{n}y_{n}\rangle$ where $L_{i}=(a_{1},\ldots,a_{i-1}:a_{i})$ and $L_{1}\subseteq L_{2}\subseteq\ldots\subseteq L_{n}$ (cf.
 [17]).
@@
 From the minimal bigraded free resolution of the $\operatorname{Sym}(I)$ in Complex (13), it is clear that the $y$-regularity of the $\operatorname{Sym}(I)$ is 00.
-The result then follows from [30, Corollary 3.2].
-Further when the ideal is of linear type, $I$ is generated by a $d$-sequence can be seen as a consequence of [25, Theorem 2.9]. ∎
+The result then follows from [30, Corollary 3.2(i)].
+Further when the ideal is of linear type, from Complex (13), $y$-regularity of the $\mathcal{R}(I)$ vanishes. Then $I$ is generated by a $d$-sequence follows from [30, Corollary 3.2(ii)]. ∎
 Some important classes of ideals for which Proposition 3.1 gives the bigraded free resolutions of the corresponding symmetric algebras are the following.
@@
     $d$-sequence ideals satisfying the condition that the grade of the ideal is one more than the minimal number of generators of the ideal.
-  3. (3)
-    Almost complete intersections.
-In the remaining part of this section, we give the explicit resolutions of the Rees algebra of some classes of ideals satisfying the hypothesis of Proposition 3.1, in particular almost complete intersections of grade $2$, grade $3$ etc.
+In particular, this class includes almost complete intersections.
+In the remaining part of this section, we give the explicit resolutions of the Rees algebra of some classes of ideals satisfying the hypothesis of Theorem 3.1, in particular almost complete intersections of grade $2$, grade $3$ etc.
 3.1. Almost complete intersections of grade $2$
-Setup 2.
+Setup 3.
 Let $B$ be a standard graded polynomial ring and $I=\langle f_{1},f_{2},f_{3}\rangle\subseteq B$ be an almost complete intersection of grade $2$ and perfect with deg $f_{i}=d_{i}$, $i=1,2,3$.
-Further assume that the presentation matrix of $I$ has the property that every non-zero entry in the $i^{th}$ row is of degree $\beta_{i}$.
 Proposition 3.6.
-Consider the ideal $I$ as in Setup 2 and the notations as in Proposition 3.1. Then the minimal bigraded free resolution of the corresponding $\mathcal{R}(I)$ is of the following form:
-$0\longrightarrow\begin{matrix}S(-(d_{1}+d_{2}+d_{3}),-2)\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}S(-(d_{k}+\beta_{k}),-1)\\ \oplus\\ S(-(d_{l}+\beta_{l}),-1)\end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0$
-where $\beta_{k}$ and $\beta_{l}$, $k,l\in\mathbb{N}$, corresponds to the degrees of elements from two different columns of the presentation matrix of $I$.
+Consider the ideal $I$ as in Setup 3 and the notations as in Theorem 3.1.
+Let $\beta_{ij}$ denote the degree of the $(i,j)^{th}$ element in the presentation matrix of $I$.
+Then the minimal bigraded free resolution of the corresponding $\mathcal{R}(I)$ is of the following form:
+$0\longrightarrow\begin{matrix}S(-(d_{1}+d_{2}+d_{3}),-2)\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}S(-(d_{k}+\beta_{k1}),-1)\\ \oplus\\ S(-(d_{l}+\beta_{l2}),-1)\end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0$
+where $\beta_{k1}$ and $\beta_{l2}$ are non-zero.
 Proof.
-From Hilbert-Burch theorem [12, Theorem 20.15], the graded minimal free resolution of $B/I$ has the form,
-$0\longrightarrow\begin{matrix}B(-(d_{k}+\beta_{k}))\\ \oplus\\ B(-(d_{l}+\beta_{l}))\end{matrix}\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\begin{matrix}B(-d_{1})\\ \oplus\\ B(-d_{2})\\ \oplus\\ B(-d_{3})\end{matrix}\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}B\longrightarrow 0$
-where $\phi_{1}=\left[\begin{matrix}f_{1}&f_{2}&f_{3}\end{matrix}\right]$ and $\phi_{2}=\left[\begin{matrix}z_{11}&z_{12}\\ z_{21}&z_{22}\\ z_{31}&z_{32}\\ \end{matrix}\right]$ and further $\beta_{k}$ and $\beta_{l}$ come corresponding to some non-zero entries $z_{ks}$ and $z_{lt}$ respectively in $\phi_{2}$ such that $s-t\cong 1\mod(2)$.
+From the Hilbert-Burch theorem [13, Theorem 20.15], the graded minimal free resolution of $B/I$ has the form,
+$0\longrightarrow\begin{matrix}B(-(d_{k}+\beta_{k1}))\\ \oplus\\ B(-(d_{l}+\beta_{l2}))\end{matrix}\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\begin{matrix}B(-d_{1})\\ \oplus\\ B(-d_{2})\\ \oplus\\ B(-d_{3})\end{matrix}\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}B\longrightarrow 0$
+where $\phi_{1}=\left[\begin{matrix}f_{1}&f_{2}&f_{3}\end{matrix}\right]$ and $\phi_{2}=\left[\begin{matrix}z_{11}&z_{12}\\ z_{21}&z_{22}\\ z_{31}&z_{32}\\ \end{matrix}\right]$ and $\beta_{k1}$ and $\beta_{l2}$ come corresponding to some non-zero entries $z_{k1}$ and $z_{l2}$ respectively in $\phi_{2}$.
 With a fixed ordered basis, the Buchsbaum-Rim complex ($\mathbb{C}$-complex) has the form,
@@
 Let $z$ be a non-zero divisor in $B$ such that $f_{1}=-z(z_{21}z_{32}-z_{22}z_{31}),\,f_{2}=z(z_{11}z_{32}-z_{12}z_{31}),\text{ and }f_{3}=-z(z_{11}z_{22}-z_{12}z_{21})$, which exists from [13], then $\alpha_{1}=\begin{bmatrix}-z(z_{12}y_{1}+z_{22}y_{2}+z_{32}y_{3})\\ z(z_{11}y_{1}+z_{21}y_{2}+z_{31}y_{3})\end{bmatrix}.$
-The result thus follows from Proposition 3.1. ∎
+The result thus follows from Theorem 3.1. ∎
 Remark 3.7.
 Perfect ideals of grade $2$ are characterized by the Hilbert-Burch theorem [13].
-Thus, any $3\times 2$ matrix $X$ with entries from a graded polynomial ring $B$ satisfying the condition that grade of the ideal of the corresponding maximal minors ($I_{2}(X)$) is bounded below by $2$ and, a non-zero divisor $z$ in $B$, corresponds to an ideal $zI_{2}(X)$ which has the form mentioned in Setup 2 and thus the associated Rees algebra is resolved using Proposition 3.6.
+Thus, any $3\times 2$ matrix $X$ with entries from a graded polynomial ring $B$ that satisfies the condition that the grade of the ideal of the corresponding maximal minors $I_{2}(X)$ is bounded below by $2$, along with a non-zero divisor $z$ in $B$, corresponds to an ideal $zI_{2}(X)$, which has the form mentioned in Setup 3.
+Consequently, the associated Rees algebra is resolved using Proposition 3.6.
 3.2. Almost complete intersections of grade $3$
-Setup 3.
-Let $B=K[X]$ denote a standard graded polynomial ring with $X$ being a skew-symmetric matrix of the form,
-$X=\begin{bmatrix}0&0&0&x_{14}&x_{15}&\cdots&x_{1t}\\ 0&0&0&x_{24}&x_{25}&\cdots&x_{2t}\\ 0&0&0&x_{34}&x_{35}&\cdots&x_{3t}\\ -x_{14}&-x_{24}&-x_{34}&0&x_{45}&\cdots&x_{4t}\\ -x_{15}&-x_{25}&-x_{35}&-x_{45}&0&\cdots&x_{5t}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\cdots&\vdots\\ -x_{1t}&-x_{2t}&-x_{3t}&-x_{4t}&-x_{5t}&\cdots&0\\ \end{bmatrix}$
-of order $t$.
-Let $I=\langle f_{1},\ldots,f_{4}\rangle$ with deg $f_{i}=d_{i}$, $i=1,\ldots,4$, be an almost complete intersection ideal of $B$ of grade $3$ and type $t-3$ for $t\in\mathbb{N}$, $t\geq 5$.
 Proposition 3.8.
-Consider the ideal $I$ as in Setup 3.
+Consider the ideal $I$ as in Setup 2.
 Then the minimal bigraded free resolution of $\mathcal{R}(I)\cong S/J$ where $S=K[X,Y]$ with $Y=\begin{bmatrix}y_{1},\ldots,y_{4}\\ \end{bmatrix}$ and deg $x_{kl}=(1,0)$ and deg $y_{i}=(d_{i},1)$, $i=1,\ldots,4$ and $J$ is the defining ideal of $\mathcal{R}(I)$, has the following form:
@@
 Proof.
-Let $I$ be an almost complete intersection of grade $3$ and type $t-3$. Following can be observed from the results in [7].
+Let $I$ be an almost complete intersection of grade $3$ and type $t-3$. The following can be observed from the results in [8].
   1. (1)
@@
 where $\psi=\begin{bmatrix}f_{1}&f_{2}&f_{3}&f_{4}\\ y_{1}&y_{2}&y_{3}&y_{4}\\ \end{bmatrix}$, $\epsilon=\begin{bmatrix}f_{23}&f_{24}&f_{34}&0\\ -f_{13}&-f_{14}&0&f_{34}\\ f_{12}&0&-f_{14}&-f_{24}\\ 0&f_{12}&f_{13}&f_{23}\\ \end{bmatrix}$ and $\sigma_{2}=\begin{bmatrix}-f_{4}&-y_{4}\\ f_{3}&y_{3}\\ -f_{2}&-y_{2}\\ f_{1}&y_{1}\\ \end{bmatrix}.$
-Then the bigraded free resolutions of $\mathcal{R}(I)$ are consequences of Proposition 3.1.
+Then the bigraded free resolutions of $\mathcal{R}(I)$ are consequences of Theorem 3.1.
 Corollary 3.9.
-The bi-regularities of $\mathcal{R}(I)$ when $I$ is an almost complete intersection of grade $3$ of type $t-3$ are as follows,
+The bi-regularities of $\mathcal{R}(I)$ when $I$ is an almost complete intersection of grade $3$ of type $t-3$ in a standard graded polynomial ring are as follows,
   1. (1)
@@
     Assume $J^{\prime}=(J:f_{n+1})$ to be an equigenerated perfect ideal of grade $2$ with the degree of the generators being $d^{\prime}$, $\mu(J^{\prime})=t$, $t\in\mathbb{N}$ and $J^{\prime}$ being linearly presented.
-By Hilbert-Burch theorem, the graded minimal free resolution of $B/J^{\prime}$ has the following form:
+By the Hilbert-Burch theorem, the graded minimal free resolution of $B/J^{\prime}$ has the following form:
     $0\longrightarrow B(-(d^{\prime}+1))^{t-1}\stackrel{{\scriptstyle\zeta_{2}}}{{\longrightarrow}}B(-d^{\prime})^{t}\stackrel{{\scriptstyle\zeta_{1}}}{{\longrightarrow}}B\longrightarrow 0$
@@
     $0\longrightarrow B(-nd)\stackrel{{\scriptstyle\phi_{n}}}{{\longrightarrow}}B(-(n-1)d)^{n}\stackrel{{\scriptstyle\phi_{n-1}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\phi_{4}}}{{\longrightarrow}}\begin{matrix}B(-(d^{\prime}+d+1))^{t-1}\\ \oplus\\ B(-3d)^{{n\choose{3}}}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{3}}}{{\longrightarrow}}\begin{matrix}B(-(d^{\prime}+d))^{t}\\ \oplus\\ B(-2d)^{{n\choose 2}}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\begin{matrix}B(-d)^{n+1}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}B\longrightarrow 0$
-    As a consequence of Proposition 3.1, the following would be a (non-standard) bigraded free resolution of $\mathcal{R}(I)$,
+    As a consequence of Theorem 3.1, the following would be a (non-standard) bigraded free resolution of $\mathcal{R}(I)$,
     $0\longrightarrow\bigoplus_{l=2}^{n}S(-(n+1)d,-l)\stackrel{{\scriptstyle\delta_{n}}}{{\longrightarrow}}\begin{matrix}S(-nd,-1)\\ \oplus\\ \bigoplus_{l=2}^{n-1}S(-nd,-l)^{n+1}\end{matrix}\stackrel{{\scriptstyle\delta_{n-1}}}{{\longrightarrow}}\cdots\stackrel{{\scriptstyle\delta_{3}}}{{\longrightarrow}}\begin{matrix}S(-(d^{\prime}+d+1),-1)^{t-1}\\ \oplus\\ S(-3d,-1)^{{n\choose 3}}\\ \oplus\\ S(-3d,-2)^{n+1\choose 3}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}S(-(d^{\prime}+d),-1)^{t}\\ \oplus\\ S(-2d,-1)^{{n\choose 2}}\\ \end{matrix}\stackrel{{\scriptstyle{\delta_{1}}}}{{\longrightarrow}}S\longrightarrow 0$
@@
 Let the relations be denoted by $\{g_{1},g_{2},g_{3}\}$.
-Then they satisfy the condition that $\{g_{1},g_{2}\}$ are $S$-regular and $(\langle g_{1},g_{2}\rangle:g_{3})=\langle f_{3},y_{3}\rangle$ ([6, Lemma 3.1]), which is clearly not equigenerated.
+Then they satisfy the condition that $\{g_{1},g_{2}\}$ are $S$-regular and $(\langle g_{1},g_{2}\rangle:g_{3})=\langle f_{3},y_{3}\rangle$ ([7, Lemma 3.1]), which is clearly not equigenerated.
     2. 2)
-      A class of ideals which satisfies the condition (a) are complete intersections minimally generated by $2$ linear terms.
-For example for $I=\langle x_{1}x_{2},x_{3}x_{4},x_{2}x_{3}\rangle\subset K[x_{1},x_{2},x_{3},x_{4}]$, then $\{x_{1}x_{2},x_{3}x_{4}\}$ forms a regular sequence and $(\langle x_{1}x_{2},x_{3}x_{4}\rangle:x_{2}x_{3})=\langle x_{1},x_{4}\rangle$ is a grade $2$ perfect ideal.
+      An example of a class of ideals $J^{\prime}$ that satisfies the condition (A) are complete intersections minimally generated by $2$ linear forms.
+For example, for $I=\langle x_{1}x_{2},x_{3}x_{4},x_{2}x_{3}\rangle\subset K[x_{1},x_{2},x_{3},x_{4}]$, the set $\{x_{1}x_{2},x_{3}x_{4}\}$ forms a regular sequence, and $(\langle x_{1}x_{2},x_{3}x_{4}\rangle:x_{2}x_{3})=\langle x_{1},x_{4}\rangle$ is a grade $2$ perfect ideal.
 In general, as mentioned in Remark 3.7, all perfect ideals of grade $2$ are characterized by the Hilbert Burch theorem.
@@
     $0\longrightarrow B(-nd)\stackrel{{\scriptstyle\phi_{n}}}{{\longrightarrow}}\cdots\longrightarrow\begin{matrix}B(-(2d^{\prime}+d+1))\\ \oplus\\ B(-4d)^{{n\choose 4}}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{4}}}{{\longrightarrow}}\begin{matrix}B(-(d^{\prime}+d+1))^{t}\\ \oplus\\ B(-3d)^{{n\choose 3}}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{3}}}{{\longrightarrow}}\begin{matrix}B(-(d^{\prime}+d))^{t}\\ \oplus\\ B(-2d)^{{n\choose 2}}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{2}}}{{\longrightarrow}}\begin{matrix}B(-d)^{n+1}\\ \end{matrix}\stackrel{{\scriptstyle\phi_{1}}}{{\longrightarrow}}B\longrightarrow 0$
-    From Proposition 3.1, the following would be a bigraded free resolution of $\mathcal{R}(I)$,
+    From Theorem 3.1, the following would be a bigraded free resolution of $\mathcal{R}(I)$,
     $0\longrightarrow\bigoplus_{l=2}^{n}S(-(n+1)d,-l)\stackrel{{\scriptstyle\delta_{n}}}{{\longrightarrow}}\cdots\longrightarrow\begin{matrix}S(-(2d^{\prime}+d+1),-1)\\ \oplus\\ S(-4d,-1)^{{n\choose 4}}\\ \oplus\\ S(-4d,-2)^{n+1\choose 4}\\ \oplus\\ S(-4d,-3)^{n+1\choose 4}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{3}}}{{\longrightarrow}}\begin{matrix}S(-(d^{\prime}+d+1),-1)^{t}\\ \oplus\\ S(-3d,-1)^{{n\choose 3}}\\ \oplus\\ S(-3d,-2)^{n+1\choose 3}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}S(-(d^{\prime}+d),-1)^{t}\\ \oplus\\ S(-2d,-1)^{{n\choose 2}}\\ \end{matrix}\stackrel{{\scriptstyle{\delta_{1}}}}{{\longrightarrow}}S\longrightarrow 0$
 Note 3.11.
-In both the cases (a) and (b), $d^{\prime}\leq d$ since $J\subseteq J^{\prime}$. This needs to be kept in mind while studying the properties of diagonals in the next section.
+In both cases (A) and (B), we have $d^{\prime}\leq d$ since $J\subseteq J^{\prime}$. This should be kept in mind while studying the properties of diagonals in the next section.
 4. Application to diagonal subalgebras of almost complete intersections
@@
 […]
-Consider the notations as in Proposition 3.1.
+Consider the notations as in Theorem 3.1.
 4.1. Koszulness of $\mathcal{R}(I)_{\Delta}$
-Let $I=\langle f_{1},\ldots,f_{n+1}\rangle$ be an equigenerated ideal of $B$ satisfying the conditions in Proposition 3.1 with deg $f_{i}=d$.
+Let $I=\langle f_{1},\ldots,f_{n+1}\rangle$ be an equigenerated ideal of $B$ satisfying the conditions in Theorem 3.1 with deg $f_{i}=d$.
 Then it is possible to associate a standard bigraded structure to the corresponding Rees algebra by considering it to be the quotient of a standard bigraded polynomial ring $S$ with deg $x_{i}=(1,0)$ and deg $y_{j}=(0,1)$ for $i=1,\ldots,m$ and $j=1,\ldots,n+1$.
-From the form of the differentials mentioned in Proposition 3.1, the standard bigraded free resolution of the Rees algebra will take the following form,
+From the form of the differentials mentioned in Theorem 3.1, the standard bigraded free resolution of the Rees algebra will take the following form,
 $\cdots\stackrel{{\scriptstyle\delta_{4}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 3}S(-j+d,-1)^{\beta_{4j}}\\ \oplus\\ C_{2}^{\prime}\end{matrix}\stackrel{{\scriptstyle\delta_{3}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 2}S(-j+d,-1)^{\beta_{3j}}\\ \oplus\\ C_{1}^{\prime}\end{matrix}\stackrel{{\scriptstyle\delta_{2}}}{{\longrightarrow}}\begin{matrix}\bigoplus_{j\geq 1}S(-j+d,-1)^{\beta_{2j}}\\ \end{matrix}\stackrel{{\scriptstyle\delta_{1}}}{{\longrightarrow}}S\longrightarrow 0.$ (19)
@@
 Having defined standard bigraded structure on $\mathcal{R}(I)$, it makes sense to talk about the Koszulness of the corresponding diagonals, since $\Delta$ being an exact functor, when applied to the standard bigraded resolution of $\mathcal{R}(I)$ gives a standard graded resolution of $\mathcal{R}(I)_{\Delta}$.
-It is clear from Complex (10) that the $y$-regularity of $\mathcal{R}(I)$ vanishes, and thus $\mathcal{R}(I)_{\Delta}$ is Koszul for $e>0$ [25, Theorem 2.9].
-Thus only the $x$-shifts need to be looked into.
 Let $i$ denote the homological degree and $t_{i}^{B}(B/I)$ be as defined in Section 1.
-Then for $a_{max}^{i}=\text{max}\{t_{i+1}^{B}(B/I)+d,(i-1)d\}$, from Remark $(1)$, $\mathcal{R}(I)_{\Delta}$ is Koszul if $\dfrac{a_{max}^{i}}{c}\leq i+1$ for all $i\geq 1$.
-Case 1: $a_{max}^{i}=(i-1)d$. Then for $c\geq d$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
-Case 2: $a_{max}^{i}=t_{i+1}^{B}(B/I)-d$. Then for $c\geq\dfrac{t_{i+1}^{B}(B/I)-d}{i+1}$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
+Let $a_{\max}^{i}$ and $b_{\max}^{i}$ denote the maximum shifts in the first and second degrees, respectively, at the $i^{th}$ homological degree of the Complex (19).
+Clearly,
+$a_{\max}^{i}=\max\{t_{i+1}^{B}(B/I)+d,(i-1)d\}$
+and
+$b_{\max}^{i}=\max\{j+2\mid j=0,\ldots,i-1\}=i+1.$
+Then from Remark 1.4(1), $\mathcal{R}(I)_{\Delta}$ is Koszul if $\dfrac{a_{max}^{i}}{c}\leq i+1$ and $\dfrac{b_{\max}^{i}}{e}\leq i+1$ for all $i\geq 1$.
+Since $b_{\max}^{i}=i+1$, the condition $\dfrac{b_{\max}^{i}}{e}\leq i+1$ holds true for all $e>0$.
+It remains to check the values of $c$ for which $\dfrac{a_{\max}^{i}}{c}\leq i+1.$ For this, we need to consider the following two cases from the Complex (19),
+Case 1: $a_{max}^{i}=(i-1)d$.
+Then $\dfrac{a_{max}^{i}}{c}\leq i+1\Longleftrightarrow\dfrac{(i-1)d}{c}\leq i+1\Longleftrightarrow\dfrac{i-1}{i+1}d\leq c\Longleftrightarrow c\geq d$ since $\left\lvert{\dfrac{i-1}{i+1}}\right\rvert\leq 1$.
+Thus for $c\geq d$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
+Case 2: $a_{max}^{i}=t_{i+1}^{B}(B/I)-d$. Then for $c\geq\dfrac{t_{i+1}^{B}(B/I)-d}{i+1}$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
 Let $\gamma=\text{max}\{\dfrac{t_{i+1}^{B}(B/I)-d}{i+1}|\,1\leq i\leq\operatorname{proj\,dim}_{B}(B/I)\}$. Then for $c\geq\max\{\gamma,d\}$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
-As a consequence of the discussion above, following can be observed regarding the Koszulness of the diagonals of $\mathcal{R}(I)$ for the class of ideals discussed in Section 3.
+As a consequence of the discussion above, the following can be observed regarding the Koszulness of the diagonals of $\mathcal{R}(I)$ for the class of ideals discussed in Section 3.
 Proposition 4.1.
@@
   1. (1)
-    Consider the assumptions as in Setup 2 with $d_{1}=d_{2}=d_{3}=d$, $d>1$ and $I$ being linearly presented. Then $\mathcal{R}(I)_{\Delta}$ is Koszul for $c\geq\frac{d}{3}$ and $e>0$.
+    Consider the assumptions as in Setup 3 with $d_{1}=d_{2}=d_{3}=d$, $d>1$ and $I$ being linearly presented. Then $\mathcal{R}(I)_{\Delta}$ is Koszul for $c\geq\frac{d}{3}$ and $e>0$.
   2. (2)
-    Consider the assumptions as in Subsection 3.3$(a)$ and $(b)$. Then for $c\geq\big{(}\frac{n-1}{n}\big{)}d$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
+    Let $I$ be an equigenerated ideal of height $n$ with a generating set $\{f_{1},\ldots,f_{n+1}\}$ such that $f_{1},\ldots,f_{n}$ forms a regular sequence.
+Further, let $J=\langle f_{1},\ldots,f_{n}\rangle$ and $J^{\prime}=(J:f_{n+1})$ satisfy the assumptions as in Subsection 3.3 (A) or (B).
+Then for $c\geq\big{(}\frac{n-1}{n}\big{)}d$ and $e>0$, $\mathcal{R}(I)_{\Delta}$ is Koszul.
 Remark 4.2.
 If $I$ is an almost complete intersection ideal of grade $2$ and perfect generated in degree $1$, then $\mathcal{R}(I)_{\Delta}$ is Koszul for all $\Delta$.
-The previous proposition thus gives bounds on the diagonals of Rees algebra associated to almost complete intersections for which they are Koszul, similar to the ones given in [9, 6, 23] for Rees algebras corresponding to complete intersections.
+Note that Proposition 4.1 provides bounds on the diagonals of the Rees algebra associated with almost complete intersections, for which they are Koszul, similar to those found for the Rees algebra corresponding to complete intersections [10, 7, 24], residual intersections [1] and maximal order Pfaffians [25].
 4.2. Cohen-Macaulayness of $\mathcal{R}(I)_{\Delta}$
@@
 Proposition 4.3.
-Let $I$ be a graded ideal of $B$ satisfying the conditions in any of the subsections 3.1, 3.2 or 3.3.
+Let $I=\langle f_{1},\ldots,f_{n+1}\rangle$ be a graded ideal in a standard graded polynomial ring $B=K[x_{1},\ldots,x_{m}]$ with the degree of $f_{i}=d_{i}$, satisfying either of the following conditions:
+  1. (1)
+    $I$ be an almost complete intersection ideal satisfying the assumptions as in Setup 2 or Setup 3.
+  2. (2)
+    The height of the ideal $I$ is $n$ with $f_{1},\ldots,f_{n}$ forming a regular sequence.
+Further, let $J=\langle f_{1},\ldots,f_{n}\rangle$ and $J^{\prime}=(J:f_{n+1})$ satisfy the assumptions as in Subsection 3.3 (A) or (B).
+Set $d=\max\{d_{1},\ldots,d_{n+1}\}$ and $u=\sum_{i=1}^{n+1}d_{i}$.
 Then $\mathcal{R}(I)$ is Cohen-Macaulay and $\mathcal{R}(I)_{\Delta}$ is Cohen-Macaulay for all $e>0$ and $c>\max\{\min\{(e-1)d+u-m,e(u-m)\},\min\{(e-1)d+u-d_{1},e(u-d_{1})\},ed\}$.
@@
 Acknowledgement.
-The Core Research Grant (CRG/2023/007668) from the Science and Engineering Research Board, India, partially supports the first author.
+The Core Research Grant (CRG/2023/007668) from the Science and Engineering Research Board, ANRF, India, partially supports the first author.
 INSPIRE Fellowship, DST, India, financially supports the second author.